GNU M4, version 1.4.7

A powerful macro processor Edition 1.4.7, 23 September 2006

by René Seindal

This manual is for GNU M4 (version 1.4.7, 23 September 2006), a package containing an implementation of the m4 macro language. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1990, 1991, 1992, 1993, 1994, 2004, 2005, 2006 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License.”

Table of Contents

1 Introduction and preliminaries . . . . . . . . . . . . . . . . 3

1.1 Introduction to m4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Historical references . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Invoking m4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Problems and bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.5 Using this manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2 Lexical and syntactic conventions . . . . . . . . . . . . 11

2.1 Macro names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.2 Quoting input to m4. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.3 Comments in m4 input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.4 Other kinds of input tokens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.5 How m4 copies input to output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

3 How to invoke macros. . . . . . . . . . . . . . . . . . . . . . . . . 13

3.1 Macro invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.2 Preventing macro invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.3 Macro arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.4 On Quoting Arguments to macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.5 Macro expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

4 How to define new macros . . . . . . . . . . . . . . . . . . . . 17

4.1 Defining a macro. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.2 Arguments to macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.3 Special arguments to macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 4.4 Deleting a macro. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.5 Renaming macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.6 Temporarily redefining macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.7 Indirect call of macros. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.8 Indirect call of builtins . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

5 Conditionals, loops, and recursion . . . . . . . . . . . 27

5.1 Testing if a macro is defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 5.2 If-else construct, or multibranch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 5.3 Loops and recursion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

6 How to debug macros and input . . . . . . . . . . . . . 31

6.1 Displaying macro definitions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 6.2 Tracing macro calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 6.3 Controlling debugging output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.4 Saving debugging output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

7 Input control. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

7.1 Deleting whitespace in input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.2 Changing the quote characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.3 Changing the comment delimiters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.4 Changing the lexical structure of words. . . . . . . . . . . . . . . . . . . . . . . . 42 7.5 Saving text until end of input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

8 File inclusion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

8.1 Including named files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8.2 Searching for include files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

9 Diverting and undiverting output . . . . . . . . . . . . 49

9.1 Diverting output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 9.2 Undiverting output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 9.3 Diversion numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 9.4 Discarding diverted text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

10 Macros for text handling . . . . . . . . . . . . . . . . . . . . 55

10.1 Calculating length of strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 10.2 Searching for substrings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 10.3 Searching for regular expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 10.4 Extracting substrings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 10.5 Translating characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 10.6 Substituting text by regular expression . . . . . . . . . . . . . . . . . . . . . . . 57 10.7 Formatted output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

11 Macros for doing arithmetic. . . . . . . . . . . . . . . . . 61

11.1 Decrement and increment operators . . . . . . . . . . . . . . . . . . . . . . . . . . 61 11.2 Evaluating integer expressions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

12 Macros for running shell commands . . . . . . . . 65

12.1 Determining the platform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 12.2 Executing simple commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 12.3 Reading the output of commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 12.4 Exit status. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 12.5 Making temporary files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

13 Miscellaneous builtin macros . . . . . . . . . . . . . . . . 69

13.1 Printing error messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 13.2 Printing current location . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 13.3 Exiting from m4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

14 Fast loading of frozen state. . . . . . . . . . . . . . . . . . 73

14.1 Using frozen files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 14.2 Frozen file format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

15 Compatibility with other versions of m4 …. 77

15.1 Extensions in GNU m4. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 15.2 Facilities in System V m4 not in GNU m4. . . . . . . . . . . . . . . . . . . . . . 78 15.3 Other incompatibilities. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

16 Correct version of some examples . . . . . . . . . . 83

Appendix A How to make copies of this manual

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

A.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

Appendix B Indices of concepts and macros . . 93

B.1 Index for many concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 B.2 Index for all m4 macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 iv GNU M4 macro processor

GNU m4 is an implementation of the traditional UNIX macro processor. It is mostly SVR4 compatible, although it has some extensions (for example, handling more than 9 positional parameters to macros). m4 also has builtin functions for including files, running shell commands, doing arithmetic, etc. Autoconf needs GNU m4 for generating ‘configure’ scripts, but not for running them.

GNU m4 was originally written by René Seindal, with subsequent changes by François Pinard and other volunteers on the Internet. All names and email addresses can be found in the files ‘AUTHORS’ and ‘THANKS’ from the GNU M4 distribution.

This is release 1.4.7. It is now considered stable: future releases in the 1.4.x series are only meant to fix bugs, increase speed, or improve documentation. However…

An experimental feature, which would improve m4 usefulness, allows for changing the syntax for what is a word in m4. You should use:

     ./configure --enable-changeword

if you want this feature compiled in. The current implementation slows down m4 considerably and is hardly acceptable. In the future, m4 2.0 will come with a different set of new features that provide similar capabilities, but without the inefficiencies, so changeword will go away and you should not count on it.

2 GNU M4 macro processor

1 Introduction and preliminaries

This first chapter explains what GNU m4 is, where m4 comes from, how to read and use this documentation, how to call the m4 program, and how to report bugs about it. It concludes by giving tips for reading the remainder of the manual.

The following chapters then detail all the features of the m4 language.

1.1 Introduction to m4

m4 is a macro processor, in the sense that it copies its input to the output, expanding macros as it goes. Macros are either builtin or user-defined, and can take any number of arguments. Besides just doing macro expansion, m4 has builtin functions for including named files, running shell commands, doing integer arithmetic, manipulating text in various ways, performing recursion, etc.. . . m4 can be used either as a front-end to a compiler, or as a macro processor in its own right.

The m4 macro processor is widely available on all UNIXes, and has been standardized by POSIX. Usually, only a small percentage of users are aware of its existence. However, those who find it often become committed users. The popularity of GNU Autoconf, which requires GNU m4 for generating ‘configure’ scripts, is an incentive for many to install it, while these people will not themselves program in m4. GNU m4 is mostly compatible with the System V, Release 3 version, except for some minor differences. See Chapter 15 [Compatibility], page 77, for more details.

Some people find m4 to be fairly addictive. They first use m4 for simple problems, then take bigger and bigger challenges, learning how to write complex sets of m4 macros along the way. Once really addicted, users pursue writing of sophisticated m4 applications even to solve simple problems, devoting more time debugging their m4 scripts than doing real work. Beware that m4 may be dangerous for the health of compulsive programmers.

1.2 Historical references

GPM was an important ancestor of m4. See C. Stratchey: “A General Purpose Macro generator”, Computer Journal 8,3 (1965), pp. 225 ff. GPM is also succinctly described into David Gries classic “Compiler Construction for Digital Computers”.

The classic B. Kernighan and P.J. Plauger: “Software Tools”, Addison-Wesley, Inc. (1976) describes and implements a Unix macro-processor language, which inspired Dennis Ritchie to write m3, a macro processor for the AP-3 minicomputer.

Kernighan and Ritchie then joined forces to develop the original m4, as described in “The M4 Macro Processor”, Bell Laboratories (1977). It had only 21 builtin macros.

While GPM was more pure, m4 is meant to deal with the true intricacies of real life: macros can be recognized without being pre-announced, skipping whitespace or end-of-lines is easier, more constructs are builtin instead of derived, etc.

Originally, the Kernighan and Plauger macro-processor, and then m3, formed the engine for the Rational FORTRAN preprocessor, that is, the Ratfor equivalent of cpp. Later, m4 was used as a frontend for Ratfor, C and Cobol.

René Seindal released his implementation of m4, GNU m4, in 1990, with the aim of removing the artificial limitations in many of the traditional m4 implementations, such as maximum line length, macro size, or number of macros.

The late Professor A. Dain Samples described and implemented a further evolution in the form of M5: “User’s Guide to the M5 Macro Language: 2nd edition”, Electronic Announcement on comp.compilers newsgroup (1992).

François Pinard took over maintenance of GNU m4 in 1992, until 1994 when he released GNU m4 1.4, which was the stable release for 10 years. It was at this time that GNU Autoconf decided to require GNU m4 as its underlying engine, since all other implementations of m4 had too many limitations.

More recently, in 2004, Paul Eggert released 1.4.1 and 1.4.2 which addressed some long standing bugs in the venerable 1.4 release. Then in 2005 Gary V. Vaughan collected together the many patches to GNU m4 1.4 that were floating around the net and released 1.4.3 and 1.4.4. And in 2006, Eric Blake joined the team and prepared patches for the release of 1.4.5, 1.4.6, and 1.4.7.

Meanwhile, development has continued on new features for m4, such as dynamic module loading and additional builtins. When complete, GNU m4 2.0 will start a new series of releases.

1.3 Invoking m4

The format of the m4 command is:

     m4 [option...] [file...]

All options begin with ‘-’, or if long option names are used, with ‘--’. A long option name need not be written completely, any unambiguous prefix is sufficient. Unless POSIXLY_ CORRECT is set in the environment, options may be intermixed with files. The argument ‘--’ is a marker to denote the end of options.

With short options, options that do not take arguments may be combined into a single command line argument with subsequent options, options with mandatory arguments may be provided either as a single command line argument or as two arguments, and options with optional arguments must be provided as a single argument. In other words, without

POSIXLY_CORRECT, m4 -QPDfoo -d a -d+f is equivalent to m4 -Q -P -D foo -d -d+f -- a,

although the latter form is considered canonical. (With POSIXLY_CORRECT, it is equivalent

to m4 -Q -P -D foo -d -- a ./-d+f).

With long options, options with mandatory arguments may be provided with an equal sign (‘=’) in a single argument, or as two arguments, and options with optional arguments must be provided as a single argument. In other words, m4 --def foo --debug a is equivalent to m4 --define=foo --debug= -- a, although the latter form is considered canonical (not to mention more robust, in case a future version of m4 introduces an option named

‘--default’).

m4 understands the following options, grouped by functionality. Several options control the overall operation of m4:

--help Print a help summary on standard output, then immediately exit m4 without reading any input files.

--version

Print the version number of the program on standard output, then immediately exit m4 without reading any input files.

-E
--fatal-warnings

Stop execution and exit m4 once the first warning has been issued, considering all of them to be fatal.

-i
--interactive

-e Makes this invocation of m4 interactive. This means that all output will be unbuffered, and interrupts will be ignored. The spelling ‘-e’ exists for compatibility with other m4 implementations, and issues a warning because it may be withdrawn in a future version of GNU M4.

-P
--prefix-builtins

Internally modify all builtin macro names so they all start with the prefix ‘m4_’. For example, using this option, one should write ‘m4_define’ instead of ‘define’, and ‘m4___file__’ instead of ‘__file__’. This option has no effect if ‘-R’ is also specified.

-Q
--quiet

--silent Suppress warnings, such as missing or superfluous arguments in macro calls, or treating the empty string as zero.

-W REGEXP
--word-regexp=REGEXP

Use REGEXP as an alternative syntax for macro names. This experimental option will not be present on all GNU m4 implementations (see Section 7.4 [Changeword], page 42).

Several options allow m4 to behave more like a preprocessor. Macro definitions and deletions can be made on the command line, the search path can be altered, and the output file can track where the input came from. These features occur with the following options:

-D NAME[=VALUE]
--define=NAME[=VALUE]

This enters NAME into the symbol table, before any input files are read. If ‘=VALUE’ is missing, the value is taken to be the empty string. The VALUE can be any string, and the macro can be defined to take arguments, just as if it was defined from within the input. This option may be given more than once; order is significant, and redefining the same NAME loses the previous value.

-I DIRECTORY
--include=DIRECTORY

Make m4 search DIRECTORY for included files that are not found in the current working directory. See Section 8.2 [Search Path], page 48, for more details. This option may be given more than once.

-s
--synclines

Generate synchronization lines, for use by the C preprocessor or other similar tools. This is useful, for example, when m4 is used as a front end to a compiler.

Source file name and line number information is conveyed by directives of the form ‘#line linenum "file"’, which are inserted as needed into the middle of the output. Such directives mean that the following line originated or was expanded from the contents of input file file at line linenum. The ‘"file"’ part is often omitted when the file name did not change from the previous directive. Synchronization directives are always given on complete lines by themselves. When a synchronization discrepancy occurs in the middle of an output line, the associated synchronization directive is delayed until the beginning of the next generated line.

-U NAME
--undefine=NAME

This deletes any predefined meaning NAME might have. Obviously, only predefined macros can be deleted in this way. This option may be given more than once; undefining a NAME that does not have a definition is silently ignored.

There are some limits within m4 that can be tuned. For compatibility, m4 also accepts some options that control limits in other implementations, but which are automatically unbounded (limited only by your hardware constraints) in GNU m4.

-G
--traditional

Suppress all the extensions made in this implementation, compared to the System V version. See Chapter 15 [Compatibility], page 77, for a list of these.

-H NUM
--hashsize=NUM

Make the internal hash table for symbol lookup be NUM entries big. For better performance, the number should be prime, but this is not checked. The default is 509 entries. It should not be necessary to increase this value, unless you define an excessive number of macros.

-L NUM
--nesting-limit=NUM

Artificially limit the nesting of macro calls to NUM levels, stopping program execution if this limit is ever exceeded. When not specified, nesting is limited to 1024 levels. The precise effect of this option might be more correctly associated with textual nesting than dynamic recursion. It has been useful when some complex m4 input was generated by mechanical means. Most users would never need this option. If shown to be obtrusive, this option (which is still experimental) might well disappear. This option does not have the ability to break endless rescanning loops, since these do not necessarily consume much memory or stack space. Through clever usage of rescanning loops, one can request complex, time-consuming computations from m4 with useful results. Putting limitations in this area would break m4 power. There are many pathological cases: ‘define(`a', `a')a’ is only the simplest example (but see Chapter 15 [Compatibility], page 77). Expecting GNU m4 to detect these would be a little like expecting a compiler system to detect and diagnose endless loops: it is a quite hard problem in general, if not undecidable!

-B NUM
-S NUM

-T NUM These options are present for compatibility with System V m4, but do nothing in this implementation. They may disappear in future releases, and issue a warning to that effect.

-N NUM
--diversions=NUM

These options are present only for compatibility with previous versions of GNU m4, and were controlling the number of possible diversions which could be used at the same time. They do nothing, because there is no fixed limit anymore. They may disappear in future releases, and issue a warning to that effect.

GNU m4 comes with a feature of freezing internal state (see Chapter 14 [Frozen files], page 73). This can be used to speed up m4 execution when reusing a common initialization script.

-F FILE
--freeze-state=FILE

Once execution is finished, write out the frozen state on the specified FILE. It is conventional, but not required, for FILE to end in ‘.m4f’.

-R FILE
--reload-state=FILE

Before execution starts, recover the internal state from the specified frozen FILE. The options ‘-D’, ‘-U’, and ‘-t’ take effect after state is reloaded, but before the input files are read.

Finally, there are several options for aiding in debugging m4 scripts.

-d[FLAGS]
--debug[=FLAGS]

Set the debug-level according to the flags FLAGS. The debug-level controls the format and amount of information presented by the debugging functions. See Section 6.3 [Debug Levels], page 33, for more details on the format and meaning of FLAGS. If omitted, FLAGS defaults to ‘aeq’.

--debugfile=FILE
-o FILE
--error-output=FILE

Redirect dumpdef output, debug messages, and trace output to the named FILE. Warnings, error messages, and errprint output are still printed to standard error. If unspecified, debug output goes to standard error; if empty, debug output is discarded. See Section 6.4 [Debug Output], page 34, for more details. The spellings ‘-o’ and ‘--error-output’ are misleading and inconsistent with other GNU tools; for now they are silently accepted as synonyms of ‘--debugfile’, but in a future version of M4, using them will cause a warning to be issued.

-l NUM
--arglength=NUM

Restrict the size of the output generated by macro tracing to NUM characters per trace line. If unspecified or zero, output is unlimited. See Section 6.3 [Debug Levels], page 33, for more details.

-t NAME
--trace=NAME

This enables tracing for the macro NAME, at any point where it is defined. NAME need not be defined when this option is given. This option may be given more than once. See Section 6.2 [Trace], page 31, for more details.

The remaining arguments on the command line are taken to be input file names. If no names are present, the standard input is read. A file name of ‘-’ is taken to mean the standard input. It is conventional, but not required, for input files to end in ‘.m4’.

The input files are read in the sequence given. Standard input can be read more than once, so the file name ‘-’ may appear multiple times on the command line; this makes a difference when input is from a terminal or other special file type. It is an error if an input file ends in the middle of argument collection, a comment, or a quoted string.

If none of the input files invoked m4exit (see Section 13.3 [M4exit], page 70), the exit status of m4 will be 0 for success, 1 for general failure (such as problems with reading an input file), and 63 for version mismatch (see Section 14.1 [Using frozen files], page 73).

If you need to read a file whose name starts with a ‘-’, you can specify it as ‘./-file’, or use ‘--’ to mark the end of options.

1.4 Problems and bugs

If you have problems with GNU M4 or think you’ve found a bug, please report it. Before reporting a bug, make sure you’ve actually found a real bug. Carefully reread the documentation and see if it really says you can do what you’re trying to do. If it’s not clear whether you should be able to do something or not, report that too; it’s a bug in the documentation!

Before reporting a bug or trying to fix it yourself, try to isolate it to the smallest possible input file that reproduces the problem. Then send us the input file and the exact results m4 gave you. Also say what you expected to occur; this will help us decide whether the problem was really in the documentation.

Once you’ve got a precise problem, send e-mail to (Internet) bug-m4@gnu.org. Please include the version number of m4 you are using. You can get this information with the command ‘m4 --version’. Also provide details about the platform you are executing on.

Non-bug suggestions are always welcome as well. If you have questions about things that are unclear in the documentation or are just obscure features, please report them too.

1.5 Using this manual

This manual contains a number of examples of m4 input and output, and a simple notation is used to distinguish input, output and error messages from m4. Examples are set out from the normal text, and shown in a fixed width font, like this

     This is an example of an example!

To distinguish input from output, all output from m4 is prefixed by the string ‘⇒’, and all error messages by the string ‘ error ’. Thus

     Example of input line
     ⇒Output line from m4
      error and an error message

The sequence ‘^D’ in an example indicates the end of the input file. The majority of these examples are self-contained, and you can run them with similar results by invoking m4 -d. In fact, the testsuite that is bundled in the GNU M4 package consists of the examples in this document!

As each of the predefined macros in m4 is described, a prototype call of the macro will be shown, giving descriptive names to the arguments, e.g.,

[Composite] example (string, [count = ‘1’], [argument]. . .)

This is a sample prototype. There is not really a macro named example, but this documents that if there were, it would be a Composite macro, rather than a Builtin. It requires at least one argument, string. Remember that in m4, there must not be a space between the macro name and the opening parenthesis, unless it was intended to call the macro without any arguments. The brackets around count and argument show that these arguments are optional. If count is omitted, the macro behaves as if count were ‘1’, whereas if argument is omitted, the macro behaves as if it were the empty string. A blank argument is not the same as an omitted argument. For

     example, ‘example(`a')’, ‘example(`a',`1')’, and ‘example(`a',`1',)’ would be-

have identically with count set to ‘1’; while ‘example(`a',)’ and ‘example(`a',`')’ would explicitly pass the empty string for count. The ellipses (‘...’) show that the macro processes additional arguments after argument, rather than ignoring them.

All macro arguments in m4 are strings, but some are given special interpretation, e.g., as numbers, file names, regular expressions, etc. The documentation for each macro will state how the parameters are interpreted, and what happens if the argument cannot be parsed according to the desired interpretation. Unless specified otherwise, a parameter specified to be a number is parsed as a decimal, even if the argument has leading zeros; and parsing the empty string as a number results in 0 rather than an error, although a warning will be issued.

This document consistently writes and uses builtin, without a hyphen, as if it were an English word. This is how the builtin primitive is spelled within m4.

10 GNU M4 macro processor

2 Lexical and syntactic conventions

As m4 reads its input, it separates it into tokens. A token is either a name, a quoted string, or any single character, that is not a part of either a name or a string. Input to m4 can also contain comments. GNU m4 does not yet understand locales; all operations are byte-oriented rather than character-oriented. However, m4 is eight-bit clean, so you can use non-ascii characters in quoted strings (see Section 7.2 [Changequote], page 37), comments (see Section 7.3 [Changecom], page 40), and macro names (see Section 4.7 [Indir], page 24), with the exception of the nul character (the zero byte ‘'\0'’).

2.1 Macro names

A name is any sequence of letters, digits, and the character _(underscore), where the first character is not a digit. m4 will use the longest such sequence found in the input. If a name has a macro definition, it will be subject to macro expansion (see Chapter 3 [Macros], page 13). Names are case-sensitive.

Examples of legal names are: ‘foo’, ‘_tmp’, and ‘name01’.

2.2 Quoting input to m4

A quoted string is a sequence of characters surrounded by quote strings, defaulting to ` and ', where the nested begin and end quotes within the string are balanced. The value of a string token is the text, with one level of quotes stripped off. Thus

     `'

⇒

is the empty string, and double-quoting turns into single-quoting.

     ``quoted''
     ⇒`quoted'

The quote characters can be changed at any time, using the builtin macro changequote. See Section 7.2 [Changequote], page 37, for more information.

2.3 Comments in m4 input

Comments in m4 are normally delimited by the characters ‘#’ and newline. All characters between the comment delimiters are ignored, but the entire comment (including the delimiters) is passed through to the output—comments are not discarded by m4.

Comments cannot be nested, so the first newline after a ‘#’ ends the comment. The commenting effect of the begin-comment string can be inhibited by quoting it.

     `quoted text' # `commented text'
     ⇒quoted text # `commented text'
     `quoting inhibits' `#' `comments'
     ⇒quoting inhibits # comments

The comment delimiters can be changed to any string at any time, using the builtin macro changecom. See Section 7.3 [Changecom], page 40, for more information.

2.4 Other kinds of input tokens

Any character, that is neither a part of a name, nor of a quoted string, nor a comment, is a token by itself. When not in the context of macro expansion, all of these tokens are just copied to output. However, during macro expansion, whitespace characters (space, tab, newline, formfeed, carriage return, vertical tab), parentheses (‘(’ and ‘)’), comma (‘,’), and dollar (‘$’) have additional roles, explained later.

2.5 How m4 copies input to output

As m4 reads the input token by token, it will copy each token directly to the output immediately.

The exception is when it finds a word with a macro definition. In that case m4 will calculate the macro’s expansion, possibly reading more input to get the arguments. It then inserts the expansion in front of the remaining input. In other words, the resulting text from a macro call will be read and parsed into tokens again.

m4 expands a macro as soon as possible. If it finds a macro call when collecting the arguments to another, it will expand the second call first. For a running example, examine how m4 handles this input:

     format(`Result is %d', eval(`2**15'))

First, m4 sees that the token ‘format’ is a macro name, so it collects the tokens ‘(’, ‘`Result is %d'’, ‘,’, and ‘ ’, before encountering another potential macro. Sure enough, ‘eval’ is a macro name, so the nested argument collection picks up ‘(’, ‘`2**15'’, and ‘)’, invoking the eval macro with the lone argument of ‘2**15’. The expansion of ‘eval(2**15)’ is ‘32768’, which is then rescanned as the five tokens ‘3’, ‘2’, ‘7’, ‘6’, and ‘8’; and combined with the next ‘)’, the format macro now has all its arguments, as if the user had typed:

     format(`Result is %d', 32768)

The format macro expands to ‘Result is 32768’, and we have another round of scanning for the tokens ‘Result’, ‘ ’, ‘is’, ‘ ’, ‘3’, ‘2’, ‘7’, ‘6’, and ‘8’. None of these are macros, so the final output is

     ⇒Result is 32768

The order in which m4 expands the macros can be explored using the Section 6.2 [Trace], page 31 facilities of GNU m4.

This process continues until there are no more macro calls to expand and all the input has been consumed.

3 How to invoke macros

This chapter covers macro invocation, macro arguments and how macro expansion is treated.

3.1 Macro invocation

Macro invocations has one of the forms

     name

which is a macro invocation without any arguments, or

     name(arg1, arg2, ..., argn)

which is a macro invocation with n arguments. Macros can have any number of arguments. All arguments are strings, but different macros might interpret the arguments in different ways.

The opening parenthesis must follow the name directly, with no spaces in between. If it does not, the macro is called with no arguments at all.

For a macro call to have no arguments, the parentheses must be left out. The macro call

     name()

is a macro call with one argument, which is the empty string, not a call with no arguments.

3.2 Preventing macro invocation

An innovation of the m4 language, compared to some of its predecessors (like Stratchey’s GPM, for example), is the ability to recognize macro calls without resorting to any special, prefixed invocation character. While generally useful, this feature might sometimes be the source of spurious, unwanted macro calls. So, GNU m4 offers several mechanisms or techniques for inhibiting the recognition of names as macro calls.

First of all, many builtin macros cannot meaningfully be called without arguments. For any of these macros, whenever an opening parenthesis does not immediately follow their name, the builtin macro call is not triggered. This solves the most usual cases, like for ‘include’ or ‘eval’. Later in this document, the sentence “This macro is recognized only with parameters” refers to this specific provision.

There is also a command line option (‘--prefix-builtins’, or ‘-P’, see Section 1.3 [Invoking m4], page 4) that renames all builtin macro with a prefix of ‘m4_’ at startup. The option has no effect whatsoever on user defined macros. For example, with this option, one has to write m4_dnl and even m4_m4exit. It also has no effect on whether a macro requires parameters.

Another alternative is to redefine problematic macros to a name less likely to cause conflicts, See Chapter 4 [Definitions], page 17.

If your version of GNU m4 has the changeword feature compiled in, it offers far more flexibility in specifying the syntax of macro names, both builtin or user-defined. See Section 7.4 [Changeword], page 42, for more information on this experimental feature.

Of course, the simplest way to prevent a name from being interpreted as a call to an existing macro is to quote it. The remainder of this section studies a little more deeply how quoting affects macro invocation, and how quoting can be used to inhibit macro invocation.

Even if quoting is usually done over the whole macro name, it can also be done over only a few characters of this name (provided, of course, that the unquoted portions are not also a macro). It is also possible to quote the empty string, but this works only inside the name. For example:

`divert'
⇒divert
`d'ivert
⇒divert
di`ver't
⇒divert
div`'ert
⇒divert

all yield the string ‘divert’. While in both:

`'divert

⇒

divert`'

⇒ the divert builtin macro will be called, which expands to the empty string.

The output of macro evaluations is always rescanned. The following example would yield the string ‘de’, exactly as if m4 has been given ‘substr(`abcde', `3', `2')’ as input:

define(`x', `substr(ab')

⇒

define(`y', `cde, `3', `2')')

⇒

x`'y
⇒de

Unquoted strings on either side of a quoted string are subject to being recognized as macro names. In the following example, quoting the empty string allows for the second macro to be recognized as such:

define(`macro', `m')

⇒

macro(`m')macro
⇒mmacro
macro(`m')`'macro
⇒mm

Quoting may prevent recognizing as a macro name the concatenation of a macro expansion with the surrounding characters. In this example:

define(`macro', `di$1')

⇒

macro(`v')`ert'
⇒divert
macro(`v')ert

⇒ the input will produce the string ‘divert’. When the quotes were removed, the divert builtin was called instead.

3.3 Macro arguments

When a name is seen, and it has a macro definition, it will be expanded as a macro.

If the name is followed by an opening parenthesis, the arguments will be collected before the macro is called. If too few arguments are supplied, the missing arguments are taken to be the empty string. However, some builtins are documented to behave differently for a missing optional argument than for an explicit empty string. If there are too many arguments, the excess arguments are ignored. Unquoted leading whitespace is stripped off all arguments.

Normally m4 will issue warnings if a builtin macro is called with an inappropriate number of arguments, but it can be suppressed with the ‘--quiet’ command line option (or ‘--silent’, or ‘-Q’, see Section 1.3 [Invoking m4], page 4). For user defined macros, there is no check of the number of arguments given.

Macros are expanded normally during argument collection, and whatever commas, quotes and parentheses that might show up in the resulting expanded text will serve to define the arguments as well. Thus, if foo expands to ‘, b, c’, the macro call

     bar(a foo, d)

is a macro call with four arguments, which are ‘a ’, ‘b’, ‘c’ and ‘d’. To understand why the first argument contains whitespace, remember that leading unquoted whitespace is never part of an argument, but trailing whitespace always is.

It is possible for a macro’s definition to change during argument collection, in which case the expansion uses the definition that was in effect at the time the opening ‘(’ was seen.

     define(`f', `1')

⇒

     f(define(`f', `2'))
     ⇒1
     f
     ⇒2

It is an error if the end of file occurs while collecting arguments.

     define(
     ^D
      error m4:stdin:1: ERROR: end of file in argument list

3.4 On Quoting Arguments to macros

Each argument has leading unquoted whitespace removed. Within each argument, all unquoted parentheses must match. For example, if foo is a macro,

     foo(() (`(') `(')

is a macro call, with one argument, whose value is ‘() (() (’. Commas separate arguments, except when they occur inside quotes, comments, or unquoted parentheses, See Section 4.3 [Pseudo Arguments], page 19, for examples.

It is common practice to quote all arguments to macros, unless you are sure you want the arguments expanded. Thus, in the above example with the parentheses, the ‘right’ way to do it is like this:

     foo(`() (() (')

It is, however, in certain cases necessary or convenient to leave out quotes for some arguments, and there is nothing wrong in doing it. It just makes life a bit harder, if you are not careful. For consistency, this manual follows the rule of thumb that each layer of parentheses introduces another layer of single quoting, except when showing the consequences of quoting rules. This is done even when the quoted string cannot be a macro, such as with integers when you have not changed the syntax via changeword (see Section 7.4 [Changeword], page 42).

3.5 Macro expansion

When the arguments, if any, to a macro call have been collected, the macro is expanded, and the expansion text is pushed back onto the input (unquoted), and reread. The expansion text from one macro call might therefore result in more macros being called, if the calls are included, completely or partially, in the first macro calls’ expansion.

Taking a very simple example, if foo expands to ‘bar’, and bar expands to ‘Hello world’, the input

     foo

will expand first to ‘bar’, and when this is reread and expanded, into ‘Hello world’.

4 How to define new macros

Macros can be defined, redefined and deleted in several different ways. Also, it is possible to redefine a macro without losing a previous value, and bring back the original value at a later time.

4.1 Defining a macro

The normal way to define or redefine macros is to use the builtin define:

[Builtin] define (name, [expansion])

Defines name to expand to expansion. If expansion is not given, it is taken to be empty. The expansion of define is void. The macro define is recognized only with parameters.

The following example defines the macro foo to expand to the text ‘Hello World.’.

     define(`foo', `Hello world.')

⇒

     foo
     ⇒Hello world.

The empty line in the output is there because the newline is not a part of the macro definition, and it is consequently copied to the output. This can be avoided by use of the macro dnl. See Section 7.1 [Dnl], page 37, for details.

The first argument to define should be quoted; otherwise, if the macro is already defined, you will be defining a different macro. This example shows the problems with underquoting, since we did not want to redefine one:

     define(foo, one)

⇒

     define(foo, two)

⇒

     one
     ⇒two

GNU m4 normally replaces only the topmost definition of a macro if it has several definitions from pushdef (see Section 4.6 [Pushdef], page 23). Some other implementations of m4 replace all definitions of a macro with define. See Section 15.2 [Incompatibilities], page 78, for more details.

As a GNU extension, the first argument to define does not have to be a simple word. It can be any text string, even the empty string. A macro with a non-standard name cannot be invoked in the normal way, as the name is not recognized. It can only be referenced by the builtins Section 4.7 [Indir], page 24 and Section 4.5 [Defn], page 21.

Arrays and associative arrays can be simulated by using this trick.

     define(`array', `defn(format(``array[%d]'', `$1'))')

⇒

     define(`array_set', `define(format(``array[%d]'', `$1'), `$2')')

⇒

array_set(`4', `array element no. 4')

⇒

array_set(`17', `array element no. 17')

⇒

array(`4')
⇒array element no. 4
array(eval(`10 + 7'))
⇒array element no. 17

Change the %d to %s and it is an associative array.

4.2 Arguments to macros

Macros can have arguments. The nth argument is denoted by $n in the expansion text, and is replaced by the nth actual argument, when the macro is expanded. Replacement of arguments happens before rescanning, regardless of how many nesting levels of quoting appear in the expansion. Here is an example of a macro with two arguments. It simply exchanges the order of the two arguments.

define(`exch', `$2, $1')

⇒

exch(`arg1', `arg2')
⇒arg2, arg1

This can be used, for example, if you like the arguments to define to be reversed.

define(`exch', `$2, $1')

⇒

define(exch(``expansion text'', ``macro''))

⇒

macro
⇒expansion text

See Section 3.4 [Quoting Arguments], page 15, for an explanation of the double quotes. (You should try and improve this example so that clients of exch do not have to double quote. see Chapter 16 [Answers], page 83)

GNU m4 allows the number following the ‘$’ to consist of one or more digits, allowing macros to have any number of arguments. This is not so in UNIX implementations of m4, which only recognize one digit.

As a special case, the zeroth argument, $0, is always the name of the macro being expanded.

define(`test', ``Macro name: $0'')

⇒

test
⇒Macro name: test

If you want quoted text to appear as part of the expansion text, remember that quotes can be nested in quoted strings. Thus, in

define(`foo', `This is macro `foo'.')

⇒

foo

⇒This is macro foo.

The ‘foo’ in the expansion text is not expanded, since it is a quoted string, and not a name.

4.3 Special arguments to macros

There is a special notation for the number of actual arguments supplied, and for all the actual arguments.

The number of actual arguments in a macro call is denoted by $# in the expansion text. Thus, a macro to display the number of arguments given can be

define(`nargs', `$#')

⇒

nargs
⇒0
nargs()
⇒1
nargs(`arg1', `arg2', `arg3')
⇒3
nargs(`commas can be quoted, like this')
⇒1
nargs(arg1#inside comments, commas do not separate arguments
still arg1)
⇒1
nargs((unquoted parentheses, like this, group arguments))
⇒1

The notation $* can be used in the expansion text to denote all the actual arguments, unquoted, with commas in between. For example

define(`echo', `$*')

⇒

echo(arg1,    arg2, arg3 , arg4)
⇒arg1,arg2,arg3 ,arg4

Often each argument should be quoted, and the notation $@ handles that. It is just like $*, except that it quotes each argument. A simple example of that is:

define(`echo', `$@')

⇒

echo(arg1,    arg2, arg3 , arg4)
⇒arg1,arg2,arg3 ,arg4

Where did the quotes go? Of course, they were eaten, when the expanded text were reread by m4. To show the difference, try

define(`echo1', `$*')

⇒

define(`echo2', `$@')

⇒

define(`foo', `This is macro `foo'.')

⇒

echo1(foo)

⇒This is macro This is macro foo..
echo1(`foo')
⇒This is macro foo.
echo2(foo)
⇒This is macro foo.
echo2(`foo')
⇒foo

See Section 6.2 [Trace], page 31, if you do not understand this. As another example of the difference, remember that comments encountered in arguments are passed untouched to the macro, and that quoting disables comments.

define(`echo1', `$*')

⇒

define(`echo2', `$@')

⇒

define(`foo', `bar')

⇒

echo1(#foo'foo
foo)
⇒#foo'foo
⇒bar
echo2(#foo'foo
foo)
⇒#foobar
⇒bar'

A ‘$’ sign in the expansion text, that is not followed by anything m4 understands, is simply copied to the macro expansion, as any other text is.

define(`foo', `$$$ hello $$$')

⇒

foo
⇒$$$ hello $$$

If you want a macro to expand to something like ‘$12’, the judicious use of nested quoting can put a safe character between the $ and the next character, relying on the rescanning to remove the nested quote. This will prevent m4 from interpreting the $ sign as a reference to an argument.

define(`foo', `no nested quote: $1')

⇒

foo(`arg')
⇒no nested quote: arg
define(`foo', `nested quote around $: `$'1')

⇒

foo(`arg')
⇒nested quote around $: $1
define(`foo', `nested empty quote after $: $`'1')

⇒

foo(`arg')

⇒nested empty quote after $: $1
define(`foo', `nested quote around next character: $`1'')

⇒

foo(`arg')
⇒nested quote around next character: $1
define(`foo', `nested quote around both: `$1'')

⇒

foo(`arg')
⇒nested quote around both: arg

4.4 Deleting a macro

A macro definition can be removed with undefine:

[Builtin] undefine (name …)

For each argument, remove the macro name. The macro names must necessarily be quoted, since they will be expanded otherwise. The expansion of undefine is void. The macro undefine is recognized only with parameters.

foo bar blah
⇒foo bar blah
define(`foo', `some')define(`bar', `other')define(`blah', `text')

⇒

foo bar blah
⇒some other text
undefine(`foo')

⇒

foo bar blah
⇒foo other text
undefine(`bar', `blah')

⇒

foo bar blah
⇒foo bar blah

Undefining a macro inside that macro’s expansion is safe; the macro still expands to the definition that was in effect at the ‘(’.

define(`f', ``$0':$1')

⇒

f(f(f(undefine(`f')`hello world')))
⇒f:f:f:hello world
f(`bye')
⇒f(bye)

It is not an error for name to have no macro definition. In that case, undefine does nothing.

4.5 Renaming macros

It is possible to rename an already defined macro. To do this, you need the builtin defn:

[Builtin] defn (name)

Expands to the quoted definition of name. If the argument is not a defined macro, the expansion is void. If name is a user-defined macro, the quoted definition is simply the quoted expansion text. If, instead, name is a builtin, the expansion is a special token, which points to the builtin’s internal definition. This token is only meaningful as the second argument to define (and pushdef), and is silently converted to an empty string in most other contexts. The macro defn is recognized only with parameters.

Its normal use is best understood through an example, which shows how to rename

undefine to zap:
define(`zap', defn(`undefine'))

⇒

zap(`undefine')

⇒

undefine(`zap')
⇒undefine(zap)

In this way, defn can be used to copy macro definitions, and also definitions of builtin macros. Even if the original macro is removed, the other name can still be used to access the definition.

The fact that macro definitions can be transferred also explains why you should use $0, rather than retyping a macro’s name in its definition:

define(`foo', `This is `$0'')

⇒

define(`bar', defn(`foo'))

⇒

bar
⇒This is bar

Macros used as string variables should be referred through defn, to avoid unwanted expansion of the text:

define(`string', `The macro dnl is very useful
')

⇒

string
⇒The macro defn(`string')
⇒The macro dnl is very useful

⇒ However, it is important to remember that m4 rescanning is purely textual. If an un-

balanced end-quote string occurs in a macro definition, the rescan will see that embedded quote as the termination of the quoted string, and the remainder of the macro’s definition will be rescanned unquoted. Thus it is a good idea to avoid unbalanced end-quotes in macro definitions or arguments to macros.

define(`foo', a'a)

⇒

define(`a', `A')

⇒

define(`echo', `$@')

⇒

foo
⇒A'A
defn(`foo')
⇒aA'
echo(foo)
⇒AA'

Using defn to generate special tokens for builtin macros outside of expected contexts can sometimes trigger warnings. But most of the time, such tokens are silently converted to the empty string.

defn(`defn')

⇒

define(defn(`divnum'), `cannot redefine a builtin token')
 error m4:stdin:2: Warning: define: invalid macro name ignored

⇒

divnum
⇒0

4.6 Temporarily redefining macros

It is possible to redefine a macro temporarily, reverting to the previous definition at a later time. This is done with the builtins pushdef and popdef:

[Builtin] pushdef (name, [expansion])

[Builtin] popdef (name …)

Analogous to define and undefine. These macros work in a stack-like fashion. A macro is temporarily redefined with pushdef, which replaces an existing definition of name, while saving the previous definition, before the new one is installed. If there is no previous definition, pushdef behaves exactly like define. If a macro has several definitions (of which only one is accessible), the topmost definition can be removed with popdef. If there is no previous definition, popdef behaves

like undefine.

The expansion of both pushdef and popdef is void. The macros pushdef and popdef are recognized only with parameters.

define(`foo', `Expansion one.')

⇒

foo
⇒Expansion one.
pushdef(`foo', `Expansion two.')

⇒

foo
⇒Expansion two.

pushdef(`foo', `Expansion three.')

⇒

pushdef(`foo', `Expansion four.')

⇒

popdef(`foo')

⇒

foo
⇒Expansion three.
popdef(`foo', `foo')

⇒

foo
⇒Expansion one.
popdef(`foo')

⇒

foo
⇒foo

If a macro with several definitions is redefined with define, the topmost definition is replaced with the new definition. If it is removed with undefine, all the definitions are removed, and not only the topmost one.

define(`foo', `Expansion one.')

⇒

foo
⇒Expansion one.
pushdef(`foo', `Expansion two.')

⇒

foo
⇒Expansion two.
define(`foo', `Second expansion two.')

⇒

foo
⇒Second expansion two.
undefine(`foo')

⇒

foo
⇒foo

Local variables within macros are made with pushdef and popdef. At the start of the macro a new definition is pushed, within the macro it is manipulated and at the end it is popped, revealing the former definition.

It is possible to temporarily redefine a builtin with pushdef and defn.

4.7 Indirect call of macros

Any macro can be called indirectly with indir:

[Builtin] indir (name, …)

Results in a call to the macro name, which is passed the rest of the arguments. If name is not defined, an error message is printed, and the expansion is void.

The macro indir is recognized only with parameters.

This can be used to call macros with computed or “invalid” names (define allows such names to be defined):

define(`$$internal$macro', `Internal macro (name `$0')')

⇒

$$internal$macro
⇒$$internal$macro
indir(`$$internal$macro')
⇒Internal macro (name $$internal$macro)

The point is, here, that larger macro packages can have private macros defined, that will not be called by accident. They can only be called through the builtin indir.

One other point to observe is that argument collection occurs before indir invokes name, so if argument collection changes the value of name, that will be reflected in the final expansion. This is different than the behavior when invoking macros directly, where the definition that was in effect before argument collection is used.

define(`f', `1')

⇒

f(define(`f', `2'))
⇒1
indir(`f', define(`f', `3'))
⇒3
indir(`f', undefine(`f'))
 error m4:stdin:4: undefined macro `f'

⇒

4.8 Indirect call of builtins

Builtin macros can be called indirectly with builtin:

[Builtin] builtin (name, …)

Results in a call to the builtin name, which is passed the rest of the arguments. If name does not name a builtin, an error message is printed, and the expansion is void. The macro builtin is recognized only with parameters.

This can be used even if name has been given another definition that has covered the original, or been undefined so that no macro maps to the builtin.

pushdef(`define', `hidden')

⇒

undefine(`undefine')

⇒

define(`foo', `bar')
⇒hidden
foo
⇒foo
builtin(`define', `foo', `BAR')

⇒

foo
⇒BAR
undefine(`foo')
⇒undefine(foo)
foo
⇒BAR
builtin(`undefine', `foo')

⇒

foo
⇒foo

The name argument only matches the original name of the builtin, even when the ‘--prefix-builtins’ option (or ‘-P’, see Section 1.3 [Invoking m4], page 4) is in effect. This is different from indir, which only tracks current macro names.

Note that indir and builtin can be used to invoke builtins without arguments, even when they normally require parameters to be recognized; but it will provoke a warning, and result in a void expansion.

builtin
⇒builtin
builtin()
 error m4:stdin:2: undefined builtin `'

⇒

builtin(`builtin')
 error m4:stdin:3: Warning: too few arguments to builtin `builtin'

⇒

builtin(`builtin',)
 error m4:stdin:4: undefined builtin `'

⇒

5 Conditionals, loops, and recursion

Macros, expanding to plain text, perhaps with arguments, are not quite enough. We would like to have macros expand to different things, based on decisions taken at run-time. For that, we need some kind of conditionals. Also, we would like to have some kind of loop construct, so we could do something a number of times, or while some condition is true.

5.1 Testing if a macro is defined

There are two different builtin conditionals in m4. The first is ifdef:

[Builtin] ifdef (name, string-1, [string-2])

If name is defined as a macro, ifdef expands to string-1, otherwise to string-2. If string-2 is omitted, it is taken to be the empty string (according to the normal rules). The macro ifdef is recognized only with parameters.

ifdef(`foo', ``foo' is defined', ``foo' is not defined')
⇒foo is not defined
define(`foo', `')

⇒

ifdef(`foo', ``foo' is defined', ``foo' is not defined')
⇒foo is defined
ifdef(`no_such_macro', `yes', `no', `extra argument')
 error m4:stdin:4: Warning: excess arguments to builtin `ifdef' ignored
⇒no

5.2 If-else construct, or multibranch

The other conditional, ifelse, is much more powerful. It can be used as a way to introduce a long comment, as an if-else construct, or as a multibranch, depending on the number of arguments supplied:

[Builtin] ifelse (comment)
[Builtin] ifelse (string-1, string-2, equal, [not-equal])
[Builtin] ifelse (string-1, string-2, equal-1, string-3, string-4, equal-2,

…) Used with only one argument, the ifelse simply discards it and produces no output. If called with three or four arguments, ifelse expands into equal, if string-1 and string-2 are equal (character for character), otherwise it expands to not-equal. A final fifth argument is ignored, after triggering a warning. If called with six or more arguments, and string-1 and string-2 are equal, ifelse expands into equal-1, otherwise the first three arguments are discarded and the processing starts again. The macro ifelse is recognized only with parameters.

Using only one argument is a common m4 idiom for introducing a block comment, as an alternative to repeatedly using dnl. This special usage is recognized by GNU m4, so that in this case, the warning about missing arguments is never triggered.

ifelse(`some comments')

⇒

ifelse(`foo', `bar')
 error m4:stdin:2: Warning: too few arguments to builtin `ifelse'

⇒

Using three or four arguments provides decision points.

ifelse(`foo', `bar', `true')

⇒

ifelse(`foo', `foo', `true')
⇒true
define(`foo', `bar')

⇒

ifelse(foo, `bar', `true', `false')
⇒true
ifelse(foo, `foo', `true', `false')
⇒false

Notice how the first argument was used unquoted; it is common to compare the expansion of a macro with a string. With this macro, you can now reproduce the behavior of many of the builtins, where the macro is recognized only with arguments.

define(`foo', `ifelse(`$#', `0', ``$0'', `arguments:$#')')

⇒

foo
⇒foo
foo()
⇒arguments:1
foo(`a', `b', `c')
⇒arguments:3

However, ifelse can take more than four arguments. If given more than four arguments, ifelse works like a case or switch statement in traditional programming languages. If string-1 and string-2 are equal, ifelse expands into equal-1, otherwise the procedure is repeated with the first three arguments discarded. This calls for an example:

ifelse(`foo', `bar', `third', `gnu', `gnats')
 error m4:stdin:1: Warning: excess arguments to builtin `ifelse' ignored
⇒gnu
ifelse(`foo', `bar', `third', `gnu', `gnats', `sixth')

⇒

ifelse(`foo', `bar', `third', `gnu', `gnats', `sixth', `seventh')
⇒seventh
ifelse(`foo', `bar', `3', `gnu', `gnats', `6', `7', `8')
 error m4:stdin:4: Warning: excess arguments to builtin `ifelse' ignored
⇒7

Naturally, the normal case will be slightly more advanced than these examples. A common use of ifelse is in macros implementing loops of various kinds.

5.3 Loops and recursion

There is no direct support for loops in m4, but macros can be recursive. There is no limit on the number of recursion levels, other than those enforced by your hardware and operating system.

Loops can be programmed using recursion and the conditionals described previously.

There is a builtin macro, shift, which can, among other things, be used for iterating through the actual arguments to a macro:

[Builtin] shift (arg1, …)

Takes any number of arguments, and expands to all its arguments except arg1, separated by commas, with each argument quoted.

The macro shift is recognized only with parameters.

shift
⇒shift
shift(`bar')

⇒

shift(`foo', `bar', `baz')
⇒bar,baz

An example of the use of shift is this macro:

[Composite] reverse (…)

Takes any number of arguments, and reverse their order.

It is implemented as:

define(`reverse', `ifelse(`$#', `0', , `$#', `1', ``$1'',
                          `reverse(shift($@)), `$1'')')

⇒

reverse

⇒

reverse(`foo')
⇒foo
reverse(`foo', `bar', `gnats', `and gnus')
⇒and gnus, gnats, bar, foo

While not a very interesting macro, it does show how simple loops can be made with shift, ifelse and recursion. It also shows that shift is usually used with ‘$@’.

Here is an example of a loop macro that implements a simple for loop.

[Composite] forloop (iterator, start, end, text)

Takes the name in iterator, which must be a valid macro name, and successively assign it each integer value from start to end, inclusive. For each assignment to iterator, append text to the expansion of the forloop. text may refer to iterator. Any definition of iterator prior to this invocation is restored.

It can, for example, be used for simple counting:

include(`forloop.m4')

⇒

forloop(`i', `1', `8', `i ')
⇒1 2 3 4 5 6 7 8

For-loops can be nested, like:

include(`forloop.m4')

⇒

forloop(`i', `1', `4', `forloop(`j', `1', `8', ` (i, j)')
')
⇒ (1, 1) (1, 2) (1, 3) (1, 4) (1, 5) (1, 6) (1, 7) (1, 8)
⇒ (2, 1) (2, 2) (2, 3) (2, 4) (2, 5) (2, 6) (2, 7) (2, 8)
⇒ (3, 1) (3, 2) (3, 3) (3, 4) (3, 5) (3, 6) (3, 7) (3, 8)
⇒ (4, 1) (4, 2) (4, 3) (4, 4) (4, 5) (4, 6) (4, 7) (4, 8)

⇒ The implementation of the forloop macro is fairly straightforward. The forloop macro

itself is simply a wrapper, which saves the previous definition of the first argument, calls the internal macro _forloop, and re-establishes the saved definition of the first argument.

The macro _forloop expands the fourth argument once, and tests to see if it is finished. If it has not finished, it increments the iteration variable (using the predefined macro incr, see Section 11.1 [Incr], page 61), and recurses.

Here is the actual implementation of forloop, distributed as ‘examples/forloop.m4’ in this package:

undivert(`forloop.m4')
⇒divert(`-1')
⇒# forloop(var, from, to, stmt)
⇒define(`forloop',
⇒ `pushdef(`$1', `$2')_forloop(`$1', `$2', `$3', `$4')popdef(`$1')')
⇒define(`_forloop',
⇒ `$4`'ifelse($1, `$3', ,
⇒  `define(`$1', incr($1))_forloop(`$1', `$2', `$3', `$4')')')
⇒divert`'dnl

⇒ Notice the careful use of quotes. Only three macro arguments are unquoted, each for

its own reason. Try to find out why these three arguments are left unquoted, and see what happens if they are quoted.

Now, even though these two macros are useful, they are still not robust enough for general use. They lack even basic error handling of cases like start value less than final value, and the first argument not being a name. Correcting these errors are left as an exercise to the reader.

6 How to debug macros and input

When writing macros for m4, they often do not work as intended on the first try (as is the case with most programming languages). Fortunately, there is support for macro debugging in m4.

6.1 Displaying macro definitions

If you want to see what a name expands into, you can use the builtin dumpdef:

[Builtin] dumpdef (…)

Accepts any number of arguments. If called without any arguments, it displays the definitions of all known names, otherwise it displays the definitions of the names given. The output is printed to the current debug file (usually standard error), and is sorted by name. If an unknown name is encountered, a warning is printed. The expansion of dumpdef is void.

define(`foo', `Hello world.')

⇒

dumpdef(`foo')
 error foo: `Hello world.'

⇒

dumpdef(`define')
 error define: <define>

⇒ The last example shows how builtin macros definitions are displayed. The definition that

is dumped corresponds to what would occur if the macro were to be called at that point, even if other definitions are still live due to redefining a macro during argument collection.

pushdef(`f', ``$0'1')pushdef(`f', ``$0'2')

⇒

f(popdef(`f')dumpdef(`f'))
 error f: ``$0'1'
⇒f2
f(popdef(`f')dumpdef(`f'))
 error m4:stdin:3: undefined macro `f'
⇒f1

See Section 6.3 [Debug Levels], page 33, for information on controlling the details of the display.

6.2 Tracing macro calls

It is possible to trace macro calls and expansions through the builtins traceon and

traceoff:

[Builtin] traceon (…)

[Builtin] traceoff (…)

When called without any arguments, traceon and traceoff will turn tracing on and off, respectively, for all defined macros. When called with arguments, only the named macros are affected, whether or not they are currently defined.

The expansion of traceon and traceoff is void.

Whenever a traced macro is called and the arguments have been collected, the call is displayed. If the expansion of the macro call is not void, the expansion can be displayed after the call. The output is printed to the current debug file (usually standard error).

define(`foo', `Hello World.')

⇒

define(`echo', `$@')

⇒

traceon(`foo', `echo')

⇒

foo
 error m4trace: -1- foo -> `Hello World.'
⇒Hello World.
echo(`gnus', `and gnats')
 error m4trace: -1- echo(`gnus', `and gnats') -> ``gnus',`and gnats''
⇒gnus,and gnats

The number between dashes is the depth of the expansion. It is one most of the time, signifying an expansion at the outermost level, but it increases when macro arguments contain unquoted macro calls. The maximum number that will appear between dashes is controlled by the option ‘--nesting-limit’ (see Section 1.3 [Invoking m4], page 4).

Tracing by name is an attribute that is preserved whether the macro is defined or not. This allows the ‘-t’ option to select macros to trace before those macros are defined.

traceoff(`foo')

⇒

traceon(`foo')

⇒

foo
⇒foo
define(`foo', `bar')

⇒

foo
 error m4trace: -1- foo -> `bar'
⇒bar
undefine(`foo')

⇒

ifdef(`foo', `yes', `no')
⇒no
indir(`foo')
 error m4:stdin:8: undefined macro `foo'

⇒

define(`foo', `blah')

⇒

foo
 error m4trace: -1- foo -> `blah'
⇒blah

     traceoff

⇒

     foo
     ⇒blah

Tracing even works on builtins. However, defn (see Section 4.5 [Defn], page 21) does not transfer tracing status.

     traceon(`eval', `m4_divnum')

⇒

     define(`m4_eval', defn(`eval'))

⇒

     define(`m4_divnum', defn(`divnum'))

⇒

     eval(divnum)
      error m4trace: -1- eval(`0') -> `0'
     ⇒0
     m4_eval(m4_divnum)
      error m4trace: -2- m4_divnum -> `0'
     ⇒0

See Section 6.3 [Debug Levels], page 33, for information on controlling the details of the display.

6.3 Controlling debugging output

The ‘-d’ option to m4 (see Section 1.3 [Invoking m4], page 4) controls the amount of details presented, when using the macros described in the preceding sections.

The flags following the option can be one or more of the following:

a Show the actual arguments in each macro call. This applies to all macro calls if the ‘t’ flag is used, otherwise only the macros covered by calls of traceon.

c Show several trace lines for each macro call. A line is shown when the macro is seen, but before the arguments are collected; a second line when the arguments have been collected and a third line after the call has completed.

e Show the expansion of each macro call, if it is not void. This applies to all macro calls if the ‘t’ flag is used, otherwise only the macros covered by calls of

          traceon.

f Show the name of the current input file in each trace output line.

i Print a message each time the current input file is changed, giving file name and input line number.

l Show the current input line number in each trace output line.

p Print a message when a named file is found through the path search mechanism (see Section 8.2 [Search Path], page 48), giving the actual file name used.

q Quote actual arguments and macro expansions in the display with the current quotes.

t Trace all macro calls made in this invocation of m4. x Add a unique ‘macro call id’ to each line of the trace output. This is useful in

connection with the ‘c’ flag above.

V A shorthand for all of the above flags.

If no flags are specified with the ‘-d’ option, the default is ‘aeq’. The examples throughout this manual assume the default flags.

There is a builtin macro debugmode, which allows on-the-fly control of the debugging output format:

[Builtin] debugmode ([flags])

The argument flags should be a subset of the letters listed above. As special cases, if the argument starts with a ‘+’, the flags are added to the current debug flags, and if it starts with a ‘-’, they are removed. If no argument is present, all debugging flags are cleared (as if no ‘-d’ was given), and with an empty argument the flags are reset to the default of ‘aeq’. The expansion of debugmode is void.

define(`foo', `FOO')

⇒

traceon(`foo')

⇒

debugmode()

⇒

foo
 error m4trace: -1- foo -> `FOO'
⇒FOO
debugmode

⇒

foo
 error m4trace: -1- foo
⇒FOO
debugmode(`+l')

⇒

foo
 error m4trace:8: -1- foo
⇒FOO

6.4 Saving debugging output

Debug and tracing output can be redirected to files using either the ‘--debugfile’ option to m4 (see Section 1.3 [Invoking m4], page 4), or with the builtin macro debugfile:

[Builtin] debugfile ([file])

Sends all further debug and trace output to file. If file is empty, debug and trace output are discarded. If debugfile is called without any arguments, debug and trace output are sent to standard error. This does not affect warnings, error messages, or errprint output, which are always sent to standard error. If file cannot be opened, the current debug file is unchanged. The expansion of debugfile is void.

traceon(`divnum')

⇒

divnum(`extra')
 error m4:stdin:2: Warning: excess arguments to builtin `divnum' ignored
 error m4trace: -1- divnum(`extra') -> `0'
⇒0
debugfile()

⇒

divnum(`extra')
 error m4:stdin:4: Warning: excess arguments to builtin `divnum' ignored
⇒0
debugfile

⇒

divnum
 error m4trace: -1- divnum -> `0'
⇒0

36 GNU M4 macro processor

7 Input control

This chapter describes various builtin macros for controlling the input to m4.

7.1 Deleting whitespace in input

The builtin dnl stands for “Discard to Next Line”:

[Builtin] dnl

All characters, up to and including the next newline, are discarded without performing any macro expansion.

The expansion of dnl is void.

It is often used in connection with define, to remove the newline that follows the call to define. Thus

     define(`foo', `Macro `foo'.')dnl A very simple macro, indeed.
     foo
     ⇒Macro foo.

The input up to and including the next newline is discarded, as opposed to the way comments are treated (see Section 2.3 [Comments], page 11).

Usually, dnl is immediately followed by an end of line or some other whitespace. GNU m4 will produce a warning diagnostic if dnl is followed by an open parenthesis. In this case, dnl will collect and process all arguments, looking for a matching close parenthesis. All predictable side effects resulting from this collection will take place. dnl will return no output. The input following the matching close parenthesis up to and including the next newline, on whatever line containing it, will still be discarded.

     dnl(`args are ignored, but side effects occur',
     define(`foo', `like this')) while this text is ignored: undefine(`foo')
      error m4:stdin:2: Warning: excess arguments to builtin `dnl' ignored
     See how `foo' was defined, foo?
     ⇒See how foo was defined, like this?

If the end of file is encountered without a newline character, a warning is issued and dnl stops consuming input.

     define(`hi', `HI')

⇒

     m4wrap(`m4wrap(`2 hi
     ')0 hi dnl 1 hi')

⇒

     ^D
      error m4: Warning: end of file treated as newline
     ⇒0 HI 2 HI

7.2 Changing the quote characters

The default quote delimiters can be changed with the builtin changequote:

[Builtin] changequote ([start = ‘`’], [end = ‘'’])

This sets start as the new begin-quote delimiter and end as the new end-quote delimiter. If any of the arguments are missing, the default quotes (` and ') are used instead of the void arguments. The expansion of changequote is void.

changequote(`[', `]')

⇒

define([foo], [Macro [foo].])

⇒

foo
⇒Macro foo.

The quotation strings can safely contain eight-bit characters. If no single character is appropriate, start and end can be of any length.

changequote(`[[[', `]]]')

⇒

define([[[foo]]], [[[Macro [[[[[foo]]]]].]]])

⇒

foo
⇒Macro [[foo]].

Changing the quotes to the empty strings will effectively disable the quoting mechanism, leaving no way to quote text.

define(`foo', `Macro `FOO'.')

⇒

changequote(, )

⇒

foo
⇒Macro `FOO'.
`foo'
⇒`Macro `FOO'.'

There is no way in m4 to quote a string containing an unmatched begin-quote, except using changequote to change the current quotes.

If the quotes should be changed from, say, ‘[’ to ‘[[’, temporary quote characters have to be defined. To achieve this, two calls of changequote must be made, one for the temporary quotes and one for the new quotes.

Macros are recognized in preference to the begin-quote string, so if a prefix of start can be recognized as a potential macro name, the quoting mechanism is effectively disabled. Unless you use changeword (see Section 7.4 [Changeword], page 42), this means that start should not begin with a letter or ‘_’ (underscore).

define(`hi', `HI')

⇒

changequote(`q', `Q')

⇒

q hi Q hi
⇒q HI Q HI

changequote

⇒

changequote(`-', `EOF')

⇒

- hi EOF hi
⇒ hi  HI

Quotes are recognized in preference to argument collection. In particular, if start is a single ‘(’, then argument collection is effectively disabled. For portability with other implementations, it is a good idea to avoid ‘(’, ‘,’, and ‘)’ as the first character in start.

define(`echo', `$#:$@:')

⇒

define(`hi', `HI')

⇒

changequote(`(',`)')

⇒

echo(hi)
⇒0::hi
changequote

⇒

changequote(`((', `))')

⇒

echo(hi)
⇒1:HI:
echo((hi))
⇒0::hi
changequote

⇒

changequote(`,', `)')

⇒

echo(hi,hi)bye)
⇒1:HIhibye:

If end is a prefix of start, the end-quote will be recognized in preference to a nested begin-quote. In particular, changing the quotes to have the same string for start and end disables nesting of quotes. When quote nesting is disabled, it is impossible to double-quote strings across macro expansions, so using the same string is not done very often.

define(`hi', `HI')

⇒

changequote(`""', `"')

⇒

""hi"""hi"
⇒hihi
""hi" ""hi"
⇒hi hi
""hi"" "hi"
⇒hi" "HI"
changequote

⇒

`hi`hi'hi'
⇒hi`hi'hi
changequote(`"', `"')

⇒

"hi"hi"hi"
⇒hiHIhi

It is an error if the end of file occurs within a quoted string.

`dangling quote
^D
 error m4:stdin:1: ERROR: end of file in string

7.3 Changing the comment delimiters

The default comment delimiters can be changed with the builtin macro changecom:

[Builtin] changecom ([start], [end])

This sets start as the new begin-comment delimiter and end as the new end-comment delimiter. If only one argument is provided, newline becomes the new end-comment delimiter. The comment delimiters can be of any length. Omitting the first argument, or using the empty string as the first argument, disables comments. The expansion of changecom is void.

define(`comment', `COMMENT')

⇒

# A normal comment
⇒# A normal comment
changecom(`/*', `*/')

⇒

# Not a comment anymore
⇒# Not a COMMENT anymore
But: /* this is a comment now */ while this is not a comment
⇒But: /* this is a comment now */ while this is not a COMMENT

Note how comments are copied to the output, much as if they were quoted strings. If you want the text inside a comment expanded, quote the begin-comment delimiter.

Calling changecom without any arguments, or with an empty string for the first argument, disables the commenting mechanism completely. To restore the original comment start of ‘#’, you must explicitly ask for it.

define(`comment', `COMMENT')

⇒

changecom

⇒

# Not a comment anymore
⇒# Not a COMMENT anymore
changecom(`#')

⇒

# comment again

⇒# comment again

The comment strings can safely contain eight-bit characters.

Comments are recognized in preference to macros. However, this is not compatible with other implementations, where macros and even quoting takes precedence over comments, so it may change in a future release. For portability, this means that start should not begin with a letter or ‘_’ (underscore), and that neither the start-quote nor the start-comment string should be a prefix of the other.

define(`hi', `HI')

⇒

changecom(`q', `Q')

⇒

q hi Q hi
⇒q hi Q HI

Comments are recognized in preference to argument collection. In particular, if start is a single ‘(’, then argument collection is effectively disabled. For portability with other implementations, it is a good idea to avoid ‘(’, ‘,’, and ‘)’ as the first character in start.

define(`echo', `$#:$@:')

⇒

define(`hi', `HI')

⇒

changecom(`(',`)')

⇒

echo(hi)
⇒0::(hi)
changecom

⇒

changecom(`((', `))')

⇒

echo(hi)
⇒1:HI:
echo((hi))
⇒0::((hi))
changecom(`,', `)')

⇒

echo(hi,hi)bye)
⇒1:HI,hi)bye:

It is an error if the end of file occurs within a comment.

changecom(`/*', `*/')

⇒

/*dangling comment
^D
 error m4:stdin:1: ERROR: end of file in comment

7.4 Changing the lexical structure of words

The macro changeword and all associated functionality is experimental. It is only available if the ‘--enable-changeword’ option was given to configure, at GNU m4 installation time. The functionality will go away in the future, to be replaced by other new features that are more efficient at providing the same capabilities. Do not rely on it. Please direct your comments about it the same way you would do for bugs.

A file being processed by m4 is split into quoted strings, words (potential macro names) and simple tokens (any other single character). Initially a word is defined by the following regular expression:

[_a-zA-Z][_a-zA-Z0-9]*

Using changeword, you can change this regular expression:

[Optional builtin] changeword (regex)

Changes the regular expression for recognizing macro names to be regex. If regex is empty, use ‘[_a-zA-Z][_a-zA-Z0-9]*’. regex must obey the constraint that every prefix of the desired final pattern is also accepted by the regular expression. If regex contains grouping parentheses, the macro invoked is the portion that matched the first group, rather than the entire matching string. The expansion of changeword is void. The macro changeword is recognized only with parameters.

Relaxing the lexical rules of m4 might be useful (for example) if you wanted to apply translations to a file of numbers:

ifdef(`changeword', `', `errprint(` skipping: no changeword support
')m4exit(`77')')dnl
changeword(`[_a-zA-Z0-9]+')

⇒

define(`1', `0')1
⇒0

Tightening the lexical rules is less useful, because it will generally make some of the builtins unavailable. You could use it to prevent accidental call of builtins, for example:

ifdef(`changeword', `', `errprint(` skipping: no changeword support
')m4exit(`77')')dnl
define(`_indir', defn(`indir'))

⇒

changeword(`_[_a-zA-Z0-9]*')

⇒

esyscmd(`foo')
⇒esyscmd(foo)
_indir(`esyscmd', `echo hi')
⇒hi

⇒ Because m4 constructs its words a character at a time, there is a restriction on the regular

expressions that may be passed to changeword. This is that if your regular expression accepts ‘foo’, it must also accept ‘f’ and ‘fo’.

ifdef(`changeword', `', `errprint(` skipping: no changeword support
')m4exit(`77')')dnl
define(`foo
', `bar
')

⇒

dnl This example wants to recognize changeword, dnl, and `foo\n'.
dnl First, we check that our regexp will match.
regexp(`changeword', `[cd][a-z]*\|foo[
]')
⇒0
regexp(`foo
', `[cd][a-z]*\|foo[
]')
⇒0
regexp(`f', `[cd][a-z]*\|foo[
]')
⇒-1
foo
⇒foo
changeword(`[cd][a-z]*\|foo[
]')

⇒

dnl Even though `foo\n' matches, we forgot to allow `f'.
foo
⇒foo
changeword(`[cd][a-z]*\|fo*[
]?')

⇒

dnl Now we can call `foo\n'.
foo
⇒bar

changeword has another function. If the regular expression supplied contains any grouped subexpressions, then text outside the first of these is discarded before symbol lookup. So:

ifdef(`changeword', `', `errprint(` skipping: no changeword support
')m4exit(`77')')dnl
ifdef(`__unix__', ,
      `errprint(` skipping: syscmd does not have unix semantics
')m4exit(`77')')dnl
changecom(`/*', `*/')dnl
define(`foo', `bar')dnl
changeword(`#\([_a-zA-Z0-9]*\)')

⇒

#esyscmd(`echo foo \#foo')
⇒foo bar

⇒

m4 now requires a ‘#’ mark at the beginning of every macro invocation, so one can use m4 to preprocess plain text without losing various words like ‘divert’.

In m4, macro substitution is based on text, while in TEX, it is based on tokens. changeword can throw this difference into relief. For example, here is the same idea represented in TEX and m4. First, the TEX version:

     \def\a{\message{Hello}}
     \catcode`\@=0
     \catcode`\\=12
     @a
     @bye
     ⇒Hello

Then, the m4 version:

     ifdef(`changeword', `', `errprint(` skipping: no changeword support
     ')m4exit(`77')')dnl
     define(`a', `errprint(`Hello')')dnl
     changeword(`@\([_a-zA-Z0-9]*\)')

⇒

     @a
     ⇒errprint(Hello)

In the TEX example, the first line defines a macro a to print the message ‘Hello’. The second line defines ^{⟨}\underline{\overline{@}}^{⟩} to be usable instead of ^{⟨}\underline{\overline{\}}^{⟩} as an escape character. The third line defines ^{⟨}\underline{\overline{\}}^{⟩} to be a normal printing character, not an escape. The fourth line invokes the macro a. So, when TEX is run on this file, it displays the message ‘Hello’.

When the m4 example is passed through m4, it outputs ‘errprint(Hello)’. The reason for this is that TEX does lexical analysis of macro definition when the macro is defined. m4 just stores the text, postponing the lexical analysis until the macro is used.

You should note that using changeword will slow m4 down by a factor of about seven, once it is changed to something other than the default regular expression. You can invoke changeword with the empty string to restore the default word definition, and regain the parsing speed.

7.5 Saving text until end of input

It is possible to ‘save’ some text until the end of the normal input has been seen. Text can be saved, to be read again by m4 when the normal input has been exhausted. This feature is normally used to initiate cleanup actions before normal exit, e.g., deleting temporary files.

To save input text, use the builtin m4wrap:

[Builtin] m4wrap (string, …)

Stores string in a safe place, to be reread when end of input is reached. As a GNU extension, additional arguments are concatenated with a space to the string.

The expansion of m4wrap is void. The macro m4wrap is recognized only with parameters.

define(`cleanup', `This is the `cleanup' action.
')

⇒

m4wrap(`cleanup')

⇒

This is the first and last normal input line.
⇒This is the first and last normal input line.
^D
⇒This is the cleanup action.

The saved input is only reread when the end of normal input is seen, and not if m4exit is used to exit m4.

It is safe to call m4wrap from saved text, but then the order in which the saved text is reread is undefined. If m4wrap is not used recursively, the saved pieces of text are reread in the opposite order in which they were saved (LIFO—last in, first out). However, this behavior is likely to change in a future release, to match POSIX, so you should not depend on this order.

Here is an example of implementing a factorial function using m4wrap:

define(`f', `ifelse(`$1', `0', `Answer: 0!=1
', eval(`$1>1'), `0', `Answer: $2$1=eval(`$2$1')
', `m4wrap(`f(decr(`$1'), `$2$1*')')')')

⇒

f(`10')

⇒

^D
⇒Answer: 10*9*8*7*6*5*4*3*2*1=3628800

Invocations of m4wrap at the same recursion level are concatenated and rescanned as usual:

define(`aa', `AA
')

⇒

m4wrap(`a')m4wrap(`a')

⇒

^D
⇒AA

however, the transition between recursion levels behaves like an end of file condition between two input files.

m4wrap(`m4wrap(`)')len(abc')

⇒

^D
 error m4: ERROR: end of file in argument list

46 GNU M4 macro processor

8 File inclusion

m4 allows you to include named files at any point in the input.

8.1 Including named files

There are two builtin macros in m4 for including files:

[Builtin] include (file)
[Builtin] sinclude (file)

Both macros cause the file named file to be read by m4. When the end of the file is reached, input is resumed from the previous input file. The expansion of include and sinclude is therefore the contents of file. If file does not exist (or cannot be read), the expansion is void, and include will fail with an error while sinclude is silent. The empty string counts as a file that does not exist. The macros include and sinclude are recognized only with parameters.

include(`none')

⇒

 error m4:stdin:1: cannot open `none': No such file or directory
include()

⇒

 error m4:stdin:2: cannot open `': No such file or directory
sinclude(`none')

⇒

sinclude()

⇒ The rest of this section assumes that m4 is invoked with the ‘-I’ option (see Section 1.3

[Invoking m4], page 4) pointing to the ‘examples’ directory shipped as part of the GNU m4 package. The file ‘examples/incl.m4’ in the distribution contains the lines:

Include file start
foo
Include file end

Normally file inclusion is used to insert the contents of a file into the input stream. The contents of the file will be read by m4 and macro calls in the file will be expanded:

define(`foo', `FOO')

⇒

include(`incl.m4')
⇒Include file start
⇒FOO
⇒Include file end

⇒ The fact that include and sinclude expand to the contents of the file can be used to

define macros that operate on entire files. Here is an example, which defines ‘bar’ to expand to the contents of ‘incl.m4’:

     define(`bar', include(`incl.m4'))

⇒

     This is `bar':  >>bar<<
     ⇒This is bar:  >>Include file start
     ⇒foo
     ⇒Include file end
     ⇒<<

This use of include is not trivial, though, as files can contain quotes, commas, and parentheses, which can interfere with the way the m4 parser works. GNU m4 seamlessly concatenates the file contents with the next character, even if the included file ended in the middle of a comment, string, or macro call. These conditions are only treated as end of file errors if specified as input files on the command line.

8.2 Searching for include files

GNU m4 allows included files to be found in other directories than the current working directory.

If a file is not found in the current working directory, and the file name is not absolute, the file will be looked for in a specified search path. First, the directories specified with the ‘-I’ option will be searched, in the order found on the command line (see Section 1.3 [Invoking m4], page 4). Second, if the M4PATH environment variable is set, it is expected to contain a colon-separated list of directories, which will be searched in order.

If the automatic search for include-files causes trouble, the ‘p’ debug flag (see Section 6.3 [Debug Levels], page 33) can help isolate the problem.

9 Diverting and undiverting output

Diversions are a way of temporarily saving output. The output of m4 can at any time be diverted to a temporary file, and be reinserted into the output stream, undiverted, again at a later time.

Numbered diversions are counted from 0 upwards, diversion number 0 being the normal output stream. The number of simultaneous diversions is limited mainly by the memory used to describe them, because GNU m4 tries to keep diversions in memory. However, there is a limit to the overall memory usable by all diversions taken altogether (512K, currently). When this maximum is about to be exceeded, a temporary file is opened to receive the contents of the biggest diversion still in memory, freeing this memory for other diversions. So, it is theoretically possible that the number of diversions be limited by the number of available file descriptors.

9.1 Diverting output

Output is diverted using divert:

[Builtin] divert ([number = ‘0’])

The current diversion is changed to number. If number is left out or empty, it is assumed to be zero. If number cannot be parsed, the diversion is unchanged.

The expansion of divert is void.

When all the m4 input will have been processed, all existing diversions are automatically undiverted, in numerical order.

     divert(`1')
     This text is diverted.
     divert

⇒

     This text is not diverted.
     ⇒This text is not diverted.
     ^D

⇒

     ⇒This text is diverted.

Several calls of divert with the same argument do not overwrite the previous diverted text, but append to it. Diversions are printed after any wrapped text is expanded.

     define(`text', `TEXT')

⇒

     divert(`1')`diverted text.'
     divert

⇒

     m4wrap(`Wrapped text preceeds ')

⇒

     ^D
     ⇒Wrapped TEXT preceeds diverted text.

If output is diverted to a non-existent diversion, it is simply discarded. This can be used to suppress unwanted output. A common example of unwanted output is the trailing newlines after macro definitions. Here is how to avoid them.

divert(`-1')
define(`foo', `Macro `foo'.')
define(`bar', `Macro `bar'.')
divert

⇒

This is a common programming idiom in m4.

Note that divert is an English word, but also an active macro without arguments. When processing plain text, the word might appear in normal text and be unintentionally swallowed as a macro invocation. One way to avoid this is to use the ‘-P’ option to rename all builtins (see Section 1.3 [Invoking m4], page 4). Another is to write a wrapper that requires a parameter to be recognized.

We decided to divert the stream for irrigation.
⇒We decided to  the stream for irrigation.
define(`divert', `ifelse(`$#', `0', ``$0'', `builtin(`$0', $@)')')

⇒

divert(-1)
Ignored text.
divert(0)

⇒

We decided to divert the stream for irrigation.
⇒We decided to divert the stream for irrigation.

9.2 Undiverting output

Diverted text can be undiverted explicitly using the builtin undivert:

[Builtin] undivert ([number]…)

Undiverts the diversions given by the arguments, in the order given. If no arguments are supplied, all diversions are undiverted, in numerical order. As a GNU extension, if number is not numeric, treat it as a file name instead.

The expansion of undivert is void.

divert(`1')
This text is diverted.
divert

⇒

This text is not diverted.
⇒This text is not diverted.
undivert(`1')

⇒

⇒This text is diverted.

⇒

Notice the last two blank lines. One of them comes from the newline following undivert, the other from the newline that followed the divert! A diversion often starts with a blank line like this.

When diverted text is undiverted, it is not reread by m4, but rather copied directly to the current output, and it is therefore not an error to undivert into a diversion. Undiverting the empty string is the same as specifying diversion 0; in either case nothing happens since the output has already been flushed.

divert(`1')diverted text
divert

⇒

undivert()

⇒

undivert(`0')

⇒

undivert
⇒diverted text

⇒

When a diversion has been undiverted, the diverted text is discarded, and it is not possible to bring back diverted text more than once.

divert(`1')
This text is diverted first.
divert(`0')undivert(`1')dnl

⇒

⇒This text is diverted first.
undivert(`1')

⇒

divert(`1')
This text is also diverted but not appended.
divert(`0')undivert(`1')dnl

⇒

⇒This text is also diverted but not appended.

Attempts to undivert the current diversion are silently ignored. Thus, when the current diversion is not 0, the current diversion does not get rearranged among the other diversions.

divert(`1')one
divert(`2')two
divert(`3')three
divert(`2')undivert`'dnl
divert`'undivert`'dnl
⇒two
⇒one
⇒three

GNU m4 allows named files to be undiverted. Given a non-numeric argument, the contents of the file named will be copied, uninterpreted, to the current output. This complements the builtin include (see Section 8.1 [Include], page 47). To illustrate the difference, the file ‘examples/foo’ contains the word ‘bar’:

define(`bar', `BAR')

⇒

undivert(`foo')
⇒bar

⇒

include(`foo')
⇒BAR

⇒ If the file is not found (or cannot be read), an error message is issued, and the expansion

is void.

9.3 Diversion numbers

The current diversion is tracked by the builtin divnum:

[Builtin] divnum

Expands to the number of the current diversion.

Initial divnum
⇒Initial 0
divert(`1')
Diversion one: divnum
divert(`2')
Diversion two: divnum
^D

⇒

⇒Diversion one: 1

⇒

⇒Diversion two: 2

9.4 Discarding diverted text

Often it is not known, when output is diverted, whether the diverted text is actually needed. Since all non-empty diversion are brought back on the main output stream when the end of input is seen, a method of discarding a diversion is needed. If all diversions should be discarded, the easiest is to end the input to m4 with ‘divert(`-1')’ followed by an explicit

‘undivert’:
divert(`1')
Diversion one: divnum
divert(`2')
Diversion two: divnum
divert(`-1')
undivert
^D

No output is produced at all.

Clearing selected diversions can be done with the following macro:

[Composite] cleardivert ([diversion]. . .)

Discard the contents of each listed diversion.

     define(`cleardivert',
     `pushdef(`_n', divnum)divert(`-1')undivert($@)divert(_n)popdef(`_n')')

⇒ It is called just like undivert, but the effect is to clear the diversions, given by the

arguments. (This macro has a nasty bug! You should try to see if you can find it and correct it. see Chapter 16 [Answers], page 83)

54 GNU M4 macro processor

10 Macros for text handling

There are a number of builtins in m4 for manipulating text in various ways, extracting substrings, searching, substituting, and so on.

10.1 Calculating length of strings

The length of a string can be calculated by len:

[Builtin] len (string)

Expands to the length of string, as a decimal number. The macro len is recognized only with parameters.

len()
⇒0
len(`abcdef')
⇒6

10.2 Searching for substrings

Searching for substrings is done with index:

[Builtin] index (string, substring)

Expands to the index of the first occurrence of substring in string. The first character in string has index 0. If substring does not occur in string, index expands to ‘-1’. The macro index is recognized only with parameters.

index(`gnus, gnats, and armadillos', `nat')
⇒7
index(`gnus, gnats, and armadillos', `dag')
⇒-1

Omitting substring evokes a warning, but still produces output.

index(`abc')
 error m4:stdin:1: Warning: too few arguments to builtin `index'
⇒0

10.3 Searching for regular expressions

Searching for regular expressions is done with the builtin regexp:

[Builtin] regexp (string, regexp, [replacement])

Searches for regexp in string. The syntax for regular expressions is the same as in GNU Emacs. See section “Syntax of Regular Expressions” in The GNU Emacs Manual. If replacement is omitted, regexp expands to the index of the first match of regexp in string. If regexp does not match anywhere in string, it expands to -1. If replacement is supplied, and there was a match, regexp changes the expansion to this argument, with ‘\n’ substituted by the text matched by the nth parenthesized sub-expression of regexp, up to nine sub-expressions. The escape ‘\&’ is replaced by the text of the entire regular expression matched. For all other characters, ‘\’ treats the next character literally. A warning is issued if there were fewer sub-expressions than the ‘\n’ requested, or if there is a trailing ‘\’. If there was no match, regexp expands to the empty string.

The macro regexp is recognized only with parameters.

regexp(`GNUs not Unix', `\<[a-z]\w+')
⇒5
regexp(`GNUs not Unix', `\<Q\w*')
⇒-1
regexp(`GNUs not Unix', `\w\(\w+\)$', `*** \& *** \1 ***')
⇒*** Unix *** nix ***
regexp(`GNUs not Unix', `\<Q\w*', `*** \& *** \1 ***')

⇒

Here are some more examples on the handling of backslash:

regexp(`abc', `\(b\)', `\\\10\a')
⇒\b0a
regexp(`abc', `b', `\1\')
 error m4:stdin:2: Warning: sub-expression 1 not present
 error m4:stdin:2: Warning: trailing \ ignored in replacement

⇒

regexp(`abc', `\(\(d\)?\)\(c\)', `\1\2\3\4\5\6')
 error m4:stdin:3: Warning: sub-expression 4 not present
 error m4:stdin:3: Warning: sub-expression 5 not present
 error m4:stdin:3: Warning: sub-expression 6 not present
⇒c

Omitting regexp evokes a warning, but still produces output.

regexp(`abc')
 error m4:stdin:1: Warning: too few arguments to builtin `regexp'
⇒0

10.4 Extracting substrings

Substrings are extracted with substr:

[Builtin] substr (string, from, [length])

Expands to the substring of string, which starts at index from, and extends for length characters, or to the end of string, if length is omitted. The starting index of a string is always 0. The expansion is empty if there is an error parsing from or length, if from is beyond the end of string, or if length is negative.

The macro substr is recognized only with parameters.

substr(`gnus, gnats, and armadillos', `6')
⇒gnats, and armadillos
substr(`gnus, gnats, and armadillos', `6', `5')
⇒gnats

Omitting from evokes a warning, but still produces output.

substr(`abc')
 error m4:stdin:1: Warning: too few arguments to builtin `substr'
⇒abc
substr(`abc',)
 error m4:stdin:2: empty string treated as 0 in builtin `substr'
⇒abc

10.5 Translating characters

Character translation is done with translit:

[Builtin] translit (string, chars, [replacement])

Expands to string, with each character that occurs in chars translated into the character from replacement with the same index. If replacement is shorter than chars, the excess characters are deleted from the expansion. If replacement is omitted, all characters in string that are present in chars are deleted from the expansion. As a GNU extension, both chars and replacement can contain character-ranges, e.g., ‘a-z’ (meaning all lowercase letters) or ‘0-9’ (meaning all digits). To include a dash ‘-’ in chars or replacement, place it first or last. It is not an error for the last character in the range to be ‘larger’ than the first. In that case, the range runs backwards, i.e., ‘9-0’ means the string ‘9876543210’. The macro translit is recognized only with parameters.

translit(`GNUs not Unix', `A-Z')
⇒s not nix
translit(`GNUs not Unix', `a-z', `A-Z')
⇒GNUS NOT UNIX
translit(`GNUs not Unix', `A-Z', `z-a')
⇒tmfs not fnix

The first example deletes all uppercase letters, the second converts lowercase to uppercase, and the third ‘mirrors’ all uppercase letters, while converting them to lowercase. The two first cases are by far the most common.

Omitting chars evokes a warning, but still produces output.

translit(`abc')
 error m4:stdin:1: Warning: too few arguments to builtin `translit'
⇒abc

10.6 Substituting text by regular expression

Global substitution in a string is done by patsubst:

[Builtin] patsubst (string, regexp, [replacement])

Searches string for matches of regexp, and substitutes replacement for each match. The syntax for regular expressions is the same as in GNU Emacs (see Section 10.3 [Regexp], page 55). The parts of string that are not covered by any match of regexp are copied to the expansion. Whenever a match is found, the search proceeds from the end of the match, so a character from string will never be substituted twice. If regexp matches a string of zero length, the start position for the search is incremented, to avoid infinite loops. When a replacement is to be made, replacement is inserted into the expansion, with ‘\n’ substituted by the text matched by the nth parenthesized sub-expression of patsubst, for up to nine sub-expressions. The escape ‘\&’ is replaced by the text of the entire regular expression matched. For all other characters, ‘\’ treats the next character literally. A warning is issued if there were fewer sub-expressions than the ‘\n’ requested, or if there is a trailing ‘\’. The replacement argument can be omitted, in which case the text matched by regexp is deleted. The macro patsubst is recognized only with parameters.

patsubst(`GNUs not Unix', `^', `OBS: ')
⇒OBS: GNUs not Unix
patsubst(`GNUs not Unix', `\<', `OBS: ')
⇒OBS: GNUs OBS: not OBS: Unix
patsubst(`GNUs not Unix', `\w*', `(\&)')
⇒(GNUs)() (not)() (Unix)()
patsubst(`GNUs not Unix', `\w+', `(\&)')
⇒(GNUs) (not) (Unix)
patsubst(`GNUs not Unix', `[A-Z][a-z]+')
⇒GN not patsubst(`GNUs not Unix', `not', `NOT\')
 error m4:stdin:6: Warning: trailing \ ignored in replacement
⇒GNUs NOT Unix

Here is a slightly more realistic example, which capitalizes individual word or whole sentences, by substituting calls of the macros upcase and downcase into the strings.

[Composite] upcase (text) [Composite] downcase (text)

[Composite] capitalize (text)

Expand to text, but with capitalization changed: upcase changes all letters to upper case, downcase changes all letters to lower case, and capitalize changes the first character of each word to upper case and the remaining characters to lower case.

define(`upcase', `translit(`$*', `a-z', `A-Z')')dnl
define(`downcase', `translit(`$*', `A-Z', `a-z')')dnl
define(`capitalize1',
       `regexp(`$1', `^\(\w\)\(\w*\)',
               `upcase(`\1')`'downcase(`\2')')')dnl
define(`capitalize',
       `patsubst(`$1', `\w+', `capitalize1(`\&')')')dnl
capitalize(`GNUs not Unix')
⇒Gnus Not Unix

While regexp replaces the whole input with the replacement as soon as there is a match, patsubst replaces each occurrence of a match and preserves non-matching pieces:

define(`patreg',

`patsubst($@)
regexp($@)')dnl
patreg(`bar foo baz Foo', `foo\|Foo', `FOO')
⇒bar FOO baz FOO
⇒FOO
patreg(`aba abb 121', `\(.\)\(.\)\1', `\2\1\2')
⇒bab abb 212
⇒bab

Omitting regexp evokes a warning, but still produces output.

patsubst(`abc')
 error m4:stdin:1: Warning: too few arguments to builtin `patsubst'
⇒abc

10.7 Formatted output

Formatted output can be made with format:

[Builtin] format (format-string, . . .)

Works much like the C function printf. The first argument format-string can contain ‘%’ specifications which are satisfied by additional arguments, and the expansion of format is the formatted string. The macro format is recognized only with parameters.

Its use is best described by a few examples:

define(`foo', `The brown fox jumped over the lazy dog')

⇒

format(`The string "%s" uses %d characters', foo, len(foo))
⇒The string "The brown fox jumped over the lazy dog" uses 38 characters
format(`%.0f', `56789.9876')
⇒56790
len(format(`%-*X', `300', `1'))
⇒300

Using the forloop macro defined in See Section 5.3 [Loops], page 29, this example shows how format can be used to produce tabular output.

include(`forloop.m4')

⇒

forloop(`i', `1', `10', `format(`%6d squared is %10d
', i, eval(i**2))')
⇒  1 squared is          1
⇒  2 squared is          4
⇒  3 squared is          9
⇒  4 squared is         16
⇒  5 squared is         25
⇒  6 squared is         36
⇒  7 squared is         49
⇒  8 squared is         64
⇒  9 squared is         81

     ⇒  10 squared is        100

⇒ The builtin format is modeled after the ANSI C ‘printf’ function, and supports these

‘%’ specifiers: ‘c’, ‘s’, ‘d’, ‘o’, ‘x’, ‘X’, ‘u’, ‘e’, ‘E’, ‘f’, ‘F’, ‘g’, ‘G’, and ‘%’; it supports field widths and precisions, and the modifiers ‘+’, ‘-’, ‘ ’, ‘0’, ‘#’, ‘h’ and ‘l’. For more details on the functioning of printf, see the C Library Manual.

For now, unrecognized specifiers are silently ignored, but it is anticipated that a future release of GNU m4 will support more specifiers, and give warnings when problems are encountered. Likewise, escape sequences are not yet recognized.

11 Macros for doing arithmetic

Integer arithmetic is included in m4, with a C-like syntax. As convenient shorthands, there are builtins for simple increment and decrement operations.

11.1 Decrement and increment operators

Increment and decrement of integers are supported using the builtins incr and decr:

[Builtin] incr (number)

[Builtin] decr (number)

Expand to the numerical value of number, incremented or decremented, respectively, by one. Except for the empty string, the expansion is empty if number could not be parsed. The macros incr and decr are recognized only with parameters.

     incr(`4')
     ⇒5
     decr(`7')
     ⇒6
     incr()
      error m4:stdin:3: empty string treated as 0 in builtin `incr'
     ⇒1
     decr()
      error m4:stdin:4: empty string treated as 0 in builtin `decr'
     ⇒-1

11.2 Evaluating integer expressions

Integer expressions are evaluated with eval:

[Builtin] eval (expression, [radix = ‘10’], [width])

Expands to the value of expression. The expansion is empty if an error is encountered while parsing the arguments. If specified, radix and width control the format of the output. The macro eval is recognized only with parameters.

Expressions can contain the following operators, listed in order of decreasing precedence.

+ - Unary plus and minus

** Exponentiation

* / % Multiplication, division and modulo

+ - Addition and subtraction

<< >> Shift left or right

== != > >= < <=

Relational operators

! Logical negation ~ Bitwise negation

& Bitwise and

^Bitwise exclusive-or

| Bitwise or

&& Logical and

|| Logical or

All operators, except exponentiation, are left associative.

Note that some older m4 implementations use ‘^’ as an alternate operator for exponentiation, although POSIX requires the C behavior of bitwise exclusive-or. On the other hand, the precedence of ‘~’ and ‘!’ are different in GNU m4 than they are in C, matching the precedence in traditional m4 implementations. This behavior is likely to change in a future version to match POSIX, so use parentheses to force the desired precedence.

Within expression, (but not radix or width), numbers without a special prefix are decimal. A simple ‘0’ prefix introduces an octal number. ‘0x’ introduces a hexadecimal number. ‘0b’ introduces a binary number. ‘0r’ introduces a number expressed in any radix between 1 and 36: the prefix should be immediately followed by the decimal expression of the radix, a colon, then the digits making the number. For radix 1, leading zeros are ignored and all remaining digits must be ‘1’; for all other radices, the digits are ‘0’, ‘1’, ‘2’, . . .. Beyond ‘9’, the digits are ‘a’, ‘b’ … up to ‘z’. Lower and upper case letters can be used interchangeably in numbers prefixes and as number digits.

Parentheses may be used to group subexpressions whenever needed. For the relational operators, a true relation returns 1, and a false relation return 0.

Here are a few examples of use of eval.

     eval(`-3 * 5')
     ⇒-15
     eval(index(`Hello world', `llo') >= 0)
     ⇒1
     eval(`0r1:0111 + 0b100 + 0r3:12')
     ⇒12
     define(`square', `eval(`('$1`)**2')')

⇒

     square(`9')
     ⇒81
     square(square(`5')`+1')
     ⇒676
     define(`foo', `666')

⇒

     eval(`foo/6')
      error m4:stdin:8: bad expression in eval: foo/6

⇒

     eval(foo/6)
     ⇒111

As the last two lines show, eval does not handle macro names, even if they expand to a valid expression (or part of a valid expression). Therefore all macros must be expanded before they are passed to eval.

All evaluation is done with 32-bit signed integers, assuming 2’s-complement with wraparound. The shift operators are defined in GNU m4 by doing an implicit bit-wise and of the right-hand operand with 0x1f, and sign-extension with right shift.

eval(0x80000000 / -1)
⇒-2147483648
eval(0x80000000 % -1)
⇒0
eval(0x7fffffff)
⇒2147483647
incr(eval(0x7fffffff))
⇒-2147483648
eval(-4 >> 33)
⇒-2

If radix is specified, it specifies the radix to be used in the expansion. The default radix is 10; this is also the case if radix is the empty string. It is an error if the radix is outside the range of 1 through 36, inclusive. The result of eval is always taken to be signed. No radix prefix is output, and for radices greater than 10, the digits are lower case. The width argument specifies the minimum output width, excluding any negative sign. The result is zero-padded to extend the expansion to the requested width. It is an error if the width is negative. On error, the expansion of eval is empty.

eval(`666', `10')
⇒666
eval(`666', `11')
⇒556
eval(`666', `6')
⇒3030
eval(`666', `6', `10')
⇒0000003030
eval(`-666', `6', `10')
⇒-0000003030
eval(`10', `', `0')
⇒10
`0r1:'eval(`10', `1', `11')
⇒0r1:01111111111
eval(`10', `16')
⇒a

64 GNU M4 macro processor

12 Macros for running shell commands

There are a few builtin macros in m4 that allow you to run shell commands from within m4.

Note that the definition of a valid shell command is system dependent. On UNIX systems, this is the typical /bin/sh. But on other systems, such as native Windows, the shell has a different syntax of commands that it understands. Some examples in this chapter assume /bin/sh, and also demonstrate how to quit early with a known exit value if this is not the case.

12.1 Determining the platform

Sometimes it is desirable for an input file to know which platform m4 is running on. GNU m4 provides several macros that are predefined to expand to the empty string; checking for their existence will confirm platform details.

[Optional builtin] __gnu__ [Optional builtin] __os2__ [Optional builtin] os2 [Optional builtin] __unix__ [Optional builtin] unix [Optional builtin] __windows__ [Optional builtin] windows

Each of these macros is conditionally defined as needed to describe the environment of m4. If defined, each macro expands to the empty string.

When GNU extensions are in effect (that is, when you did not use the ‘-G’ option, see Section 1.3 [Invoking m4], page 4), GNU m4 will define the macro __gnu__to expand to the empty string.

     __gnu__

⇒

     ifdef(`__gnu__', `Extensions are active')
     ⇒Extensions are active

On UNIX systems, GNU m4 will define __unix__by default, or unix when the ‘-G’ option is specified.

On native Windows systems, GNU m4 will define __windows__by default, or windows when the ‘-G’ option is specified.

On OS/2 systems, GNU m4 will define __os2__by default, or os2 when the ‘-G’ option is specified.

If GNU m4 does not provide a platform macro for your system, please report that as a bug.

     define(`provided', `0')

⇒

     ifdef(`__unix__', `define(`provided', incr(provided))')

⇒

     ifdef(`__windows__', `define(`provided', incr(provided))')

⇒

ifdef(`__os2__', `define(`provided', incr(provided))')

⇒

provided
⇒1

12.2 Executing simple commands

Any shell command can be executed, using syscmd:

[Builtin] syscmd (shell-command)

Executes shell-command as a shell command. The expansion of syscmd is void, not the output from shell-command! Output or error messages from shell-command are not read by m4. See Section 12.3 [Esyscmd], page 66, if you need to process the command output. Prior to executing the command, m4 flushes its buffers. The default standard input, output and error of shell-command are the same as those of m4. The macro syscmd is recognized only with parameters.

define(`foo', `FOO')

⇒

syscmd(`echo foo')
⇒foo

⇒

Note how the expansion of syscmd keeps the trailing newline of the command, as well as using the newline that appeared after the macro.

As an example of shell-command using the same standard input as m4, the command line echo "m4wrap(\`syscmd(\`cat')')" | m4 will tell m4 to read all of its input before executing the wrapped text, then hand a valid (albeit emptied) pipe as standard input for the cat subcommand. Therefore, you should be careful when using standard input (either by specifying no files, or by passing ‘-’ as a file name on the command line, see Section 1.3 [Invoking m4], page 4), and also invoking subcommands via syscmd or esyscmd that consume data from standard input. When standard input is a seekable file, the subprocess will pick up with the next character not yet processed by m4; when it is a pipe or other non-seekable file, there is no guarantee how much data will already be buffered by m4 and thus unavailable to the child.

12.3 Reading the output of commands

If you want m4 to read the output of a shell command, use esyscmd:

[Builtin] esyscmd (shell-command)

Expands to the standard output of the shell command shell-command. Prior to executing the command, m4 flushes its buffers. The default standard input and error output of shell-command are the same as those of m4. The error output of shell-command is not a part of the expansion: it will appear along with the error output of m4. The macro esyscmd is recognized only with parameters.

define(`foo', `FOO')

⇒

esyscmd(`echo foo')
⇒FOO

⇒

Note how the expansion of esyscmd keeps the trailing newline of the command, as well as using the newline that appeared after the macro.

Just as with syscmd, care must be exercised when sharing standard input between m4 and the child process of esyscmd.

12.4 Exit status

To see whether a shell command succeeded, use sysval:

[Builtin] sysval

Expands to the exit status of the last shell command run with syscmd or esyscmd. Expands to 0 if no command has been run yet.

syscmd(`false')

⇒

ifelse(sysval, `0', `zero', `non-zero')
⇒non-zero
syscmd(`exit 2')

⇒

sysval
⇒2
syscmd(`true')

⇒

sysval
⇒0
esyscmd(`false')

⇒

ifelse(sysval, `0', `zero', `non-zero')
⇒non-zero
esyscmd(`exit 2')

⇒

sysval
⇒2
esyscmd(`true')

⇒

sysval
⇒0

sysval results in 127 if there was a problem executing the command, for example, if the system-imposed argument length is exceeded, or if there were not enough resources to fork. It is not possible to distinguish between failed execution and successful execution that had an exit status of 127.

On UNIX platforms, where it is possible to detect when command execution is terminated by a signal, rather than a normal exit, the result is the signal number shifted left by eight bits.

dnl This test assumes kill is a shell builtin, and that signals are
dnl recognizable.
ifdef(`__unix__', ,
      `errprint(` skipping: syscmd does not have unix semantics
')m4exit(`77')')dnl
syscmd(`kill -13 $$')

⇒

sysval
⇒3328
esyscmd(`kill -9 $$')

⇒

sysval
⇒2304

12.5 Making temporary files

Commands specified to syscmd or esyscmd might need a temporary file, for output or for some other purpose. There is a builtin macro, maketemp, for making temporary file names:

[Builtin] maketemp (template)

Expands to a name of a new, empty file, made from the string template, which should end with the string ‘XXXXXX’. The six X characters are then replaced with random data, in order to make the file name unique. The macro maketemp is recognized only with parameters.

maketemp(`/tmp/fooXXXXXX')
⇒/tmp/fooa07346

Traditional implementations of m4 replaced the trailing ‘X’ sequence with the process id, without creating the file; meaning you only get one result no matter how many times you use maketemp on the same string. As of this release, POSIX is considering the addition of a new macro mkstemp that behaves like GNU maketemp, so a future version of GNU M4 may have changes in this area.

13 Miscellaneous builtin macros

This chapter describes various builtins, that do not really belong in any of the previous chapters.

13.1 Printing error messages

You can print error messages using errprint:

[Builtin] errprint (message, . . .)

Prints message and the rest of the arguments on the standard error output, separated by spaces. The expansion of errprint is void. The macro errprint is recognized only with parameters.

     errprint(`Invalid arguments to forloop
     ')
      error Invalid arguments to forloop

⇒

A trailing newline is not printed automatically, so it must be supplied as part of the argument, as in the example. BSD implementations of m4 do append a trailing newline on each errprint call, while some other implementations only print the first argument.

13.2 Printing current location

To make it possible to specify the location of an error, three utility builtins exist:

[Builtin] __file__

[Builtin] __line__

[Builtin] __program__

Expand to the quoted name of the current input file, the current input line number in that file, and the quoted name of the current invocation of m4.

     errprint(__program__:__file__:__line__: `input error
     ')
      error m4:stdin:1: input error

⇒

Line numbers start at 1 for each file. If the file was found due to the ‘-I’ option or M4PATH environment variable, that is reflected in the file name. The syncline option (‘-s’, see Section 1.3 [Invoking m4], page 4), and the ‘f’ and ‘l’ flags of debugmode (see Section 6.3 [Debug Levels], page 33), also use this notion of current file and line. Redefining the three location macros has no effect on syncline, debug, or warning message output. Assume this example is run in the ‘checks’ directory of the GNU M4 package, using ‘--include=../examples’ in the command line to find the file ‘incl.m4’ mentioned earlier:

     define(`foo', ``$0' called at __file__:__line__')

⇒

     foo
     ⇒foo called at stdin:2

include(`incl.m4')
⇒Include file start
⇒foo called at ../examples/incl.m4:2
⇒Include file end

⇒ Currently, all text wrapped with m4wrap (see Section 7.5 [M4wrap], page 44) behaves

as though it came from line 0 of the file “”. It is hoped that a future release of m4 can overcome this limitation and remember which file invoked the call to m4wrap.

The __program__macro behaves like ‘$0’ in shell terminology. If you invoke m4 through an absolute path or a link with a different spelling, rather than by relying on a PATH search for plain ‘m4’, it will affect how __program__expands. The intent is that you can use it to produce error messages with the same formatting that m4 produces internally. It can also be used within syscmd (see Section 12.2 [Syscmd], page 66) to pick the same version of m4 that is currently running, rather than whatever version of m4 happens to be first in PATH.

13.3 Exiting from m4

If you need to exit from m4 before the entire input has been read, you can use m4exit:

[Builtin] m4exit ([code = ‘0’])

Causes m4 to exit, with exit status code. If code is left out, the exit status is zero. If code cannot be parsed, or is outside the range of 0 to 255, the exit status is one. No further input is read, and all wrapped and diverted text is discarded.

A common use of this is to abort processing:

[Composite] fatal_error (message)

Abort processing with an error message and non-zero status. Prefix message with details about where the error occurred, and print the resulting string to standard error.

define(`fatal_error',
       `errprint(__program__:__file__:__line__`: fatal error: $*
')m4exit(`1')')

⇒

fatal_error(`this is a BAD one, buster')
 error m4:stdin:4: fatal error: this is a BAD one, buster

After this macro call, m4 will exit with exit status 1. This macro is only intended for error exits, since the normal exit procedures are not followed, e.g., diverted text is not undiverted, and saved text (see Section 7.5 [M4wrap], page 44) is not reread. (This macro has a subtle bug, when invoked from wrapped text. You should try to see if you can find it and correct it. see Chapter 16 [Answers], page 83)

m4wrap(`This text is lost to `m4exit'.')

⇒

divert(`1') And so is this.
divert

⇒

m4exit

Note that it is still possible for the exit status to be different than what was requested by m4exit. If m4 detects some other error, such as a write error on standard out, the exit status will be non-zero even if m4exit requested zero.

If standard input is seekable, then the file will be positioned at the next unread character. If it is a pipe or other non-seekable file, then there are no guarantees how much data m4 might have read into buffers, and thus discarded.

72 GNU M4 macro processor

14 Fast loading of frozen state

Some bigger m4 applications may be built over a common base containing hundreds of definitions and other costly initializations. Usually, the common base is kept in one or more declarative files, which files are listed on each m4 invocation prior to the user’s input file, or else each input file uses include.

Reading the common base of a big application, over and over again, may be time consuming. GNU m4 offers some machinery to speed up the start of an application using lengthy common bases.

14.1 Using frozen files

Suppose a user has a library of m4 initializations in ‘base.m4’, which is then used with multiple input files:

     m4 base.m4 input1.m4
     m4 base.m4 input2.m4
     m4 base.m4 input3.m4

Rather than spending time parsing the fixed contents of ‘base.m4’ every time, the user might rather execute:

     m4 -F base.m4f base.m4

once, and further execute, as often as needed:

     m4 -R base.m4f input1.m4
     m4 -R base.m4f input2.m4
     m4 -R base.m4f input3.m4

with the varying input. The first call, containing the ‘-F’ option, only reads and executes file ‘base.m4’, defining various application macros and computing other initializations. Once the input file ‘base.m4’ has been completely processed, GNU m4 produces on ‘base.m4f’ a frozen file, that is, a file which contains a kind of snapshot of the m4 internal state.

Later calls, containing the ‘-R’ option, are able to reload the internal state of m4, from ‘base.m4f’, prior to reading any other input files. This means instead of starting with a virgin copy of m4, input will be read after having effectively recovered the effect of a prior run. In our example, the effect is the same as if file ‘base.m4’ has been read anew. However, this effect is achieved a lot faster.

Only one frozen file may be created or read in any one m4 invocation. It is not possible to recover two frozen files at once. However, frozen files may be updated incrementally, through using ‘-R’ and ‘-F’ options simultaneously. For example, if some care is taken, the command:

     m4 file1.m4 file2.m4 file3.m4 file4.m4

could be broken down in the following sequence, accumulating the same output:

     m4 -F file1.m4f file1.m4
     m4 -R file1.m4f -F file2.m4f file2.m4
     m4 -R file2.m4f -F file3.m4f file3.m4
     m4 -R file3.m4f file4.m4

Some care is necessary because not every effort has been made for this to work in all cases. In particular, the trace attribute of macros is not handled, nor the current setting of changeword. Currently, m4wrap and sysval also have problems. Also, interactions for some options of m4, being used in one call and not in the next, have not been fully analyzed yet. On the other end, you may be confident that stacks of pushdef definitions are handled correctly, as well as undefined or renamed builtins, and changed strings for quotes or comments. And future releases of GNU M4 will improve on the utility of frozen files.

When an m4 run is to be frozen, the automatic undiversion which takes place at end of execution is inhibited. Instead, all positively numbered diversions are saved into the frozen file. The active diversion number is also transmitted.

A frozen file to be reloaded need not reside in the current directory. It is looked up the same way as an include file (see Section 8.2 [Search Path], page 48).

If the frozen file was generated with a newer version of m4, and contains directives that an older m4 cannot parse, attempting to load the frozen file with option ‘-R’ will cause m4 to exit with status 63 to indicate version mismatch.

14.2 Frozen file format

Frozen files are sharable across architectures. It is safe to write a frozen file on one machine and read it on another, given that the second machine uses the same or newer version of GNU m4. It is conventional, but not required, to give a frozen file the suffix of .m4f.

These are simple (editable) text files, made up of directives, each starting with a capital letter and ending with a newline (^{⟨}\underline{\overline{NL}}^{⟩}). Wherever a directive is expected, the character # introduces a comment line; empty lines are also ignored if they are not part of an embedded string. In the following descriptions, each len refers to the length of the corresponding strings str in the next line of input. Numbers are always expressed in decimal. There are no escape characters. The directives are:

C len1 , len2 ^{⟨}\underline{\overline{NL}}^{⟩} str1 str2 ^{⟨}\underline{\overline{NL}}^{⟩}

Uses str1 and str2 as the begin-comment and end-comment strings. If omitted, then ‘#’ and ^{⟨}\underline{\overline{NL}}^{⟩} are the comment delimiters.

D number, len ^{⟨}\underline{\overline{NL}}^{⟩} str ^{⟨}\underline{\overline{NL}}^{⟩}

Selects diversion number, making it current, then copy str in the current diversion. number may be a negative number for a non-existing diversion. To merely specify an active selection, use this command with an empty str. With 0 as the diversion number, str will be issued on standard output at reload time. GNU m4 will not produce the ‘D’ directive with non-zero length for diversion 0, but this can be done with manual edits. This directive may appear more than once for the same diversion, in which case the diversion is the concatenation of the various uses. If omitted, then diversion 0 is current.

F len1 , len2 ^{⟨}\underline{\overline{NL}}^{⟩} str1 str2 ^{⟨}\underline{\overline{NL}}^{⟩}

Defines, through pushdef, a definition for str1 expanding to the function whose builtin name is str2. If the builtin does not exist (for example, if the frozen file was produced by a copy of m4 compiled with changeword support, but the version of m4 reloading was compiled without it), the reload is silent, but any subsequent use of the definition of str1 will result in a warning. This directive may appear more than once for the same name, and its order, along with ‘T’, is important. If omitted, you will have no access to any builtins.

Q len1 , len2 ^{⟨}\underline{\overline{NL}}^{⟩} str1 str2 ^{⟨}\underline{\overline{NL}}^{⟩}

Uses str1 and str2 as the begin-quote and end-quote strings. If omitted, then ‘`’ and ‘'’ are the quote delimiters.

T len1 , len2 ^{⟨}\underline{\overline{NL}}^{⟩} str1 str2 ^{⟨}\underline{\overline{NL}}^{⟩}

Defines, though pushdef, a definition for str1 expanding to the text given by str2. This directive may appear more than once for the same name, and its order, along with ‘F’, is important.

V number ^{⟨}\underline{\overline{NL}}^{⟩}

Confirms the format of the file. m4 1.4.7 only creates and understands frozen files where number is 1. This directive must be the first non-comment in the file, and may not appear more than once.

76 GNU M4 macro processor

15 Compatibility with other versions of m4

This chapter describes the differences between this implementation of m4, and the implementation found under UNIX, notably System V, Release 3.

There are also differences in BSD flavors of m4. No attempt is made to summarize these here.

15.1 Extensions in GNU m4

This version of m4 contains a few facilities that do not exist in System V m4. These extra facilities are all suppressed by using the ‘-G’ command line option (see Section 1.3 [Invoking m4], page 4), unless overridden by other command line options.

• In the $n notation for macro arguments, n can contain several digits, while the System V m4 only accepts one digit. This allows macros in GNU m4 to take any number of arguments, and not only nine (see Section 4.2 [Arguments], page 18). This means that define(`foo', `$11') is ambiguous between implementations. To portably choose between grabbing the first parameter and appending 1 to the expansion, or grabbing the eleventh parameter, you can do the following:

     define(`a1', `A1')

⇒

     dnl First argument, concatenated with 1
     define(`_1', `$1')define(`first1', `_1($@)1')

⇒

     dnl Eleventh argument, portable
     define(`_9', `$9')define(`eleventh', `_9(shift(shift($@)))')

⇒

     dnl Eleventh argument, GNU style
     define(`Eleventh', `$11')

⇒

     first1(`a', `b', `c', `d', `e', `f', `g', `h', `i', `j', `k')
     ⇒A1
     eleventh(`a', `b', `c', `d', `e', `f', `g', `h', `i', `j', `k')
     ⇒k
     Eleventh(`a', `b', `c', `d', `e', `f', `g', `h', `i', `j', `k')
     ⇒k

• The divert (see Section 9.1 [Divert], page 49) macro can manage more than 9 diversions. GNU m4 treats all positive numbers as valid diversions, rather than discarding diversions greater than 9.

• Files included with include and sinclude are sought in a user specified search path, if they are not found in the working directory. The search path is specified by the ‘-I’ option and the M4PATH environment variable (see Section 8.2 [Search Path], page 48).

• Arguments to undivert can be non-numeric, in which case the named file will be included uninterpreted in the output (see Section 9.2 [Undivert], page 50).

• Formatted output is supported through the format builtin, which is modeled after the C library function printf (see Section 10.7 [Format], page 59).

• Searches and text substitution through regular expressions are supported by the regexp (see Section 10.3 [Regexp], page 55) and patsubst (see Section 10.6 [Patsubst], page 57) builtins.

• The output of shell commands can be read into m4 with esyscmd (see Section 12.3 [Esyscmd], page 66).

• There is indirect access to any builtin macro with builtin (see Section 4.8 [Builtin], page 25).

• Macros can be called indirectly through indir (see Section 4.7 [Indir], page 24).

• The name of the program, the current input file, and the current input line number are accessible through the builtins __program__, __file__, and __line__(see Section 13.2 [Location], page 69).

• The format of the output from dumpdef and macro tracing can be controlled with debugmode (see Section 6.3 [Debug Levels], page 33).

• The destination of trace and debug output can be controlled with debugfile (see Section 6.4 [Debug Output], page 34).

In addition to the above extensions, GNU m4 implements the following command line options: ‘-F’, ‘-G’, ‘-I’, ‘-L’, ‘-R’, ‘-V’, ‘-W’, ‘-d’, ‘-i’, ‘-l’, ‘--debugfile’ and ‘-t’. See Section 1.3 [Invoking m4], page 4, for a description of these options.

Also, the debugging and tracing facilities in GNU m4 are much more extensive than in most other versions of m4.

15.2 Facilities in System V m4 not in GNU m4

The version of m4 from System V contains a few facilities that have not been implemented in GNU m4 yet. Additionally, POSIX requires some behaviors that GNU m4 has not implemented yet. Relying on these behaviors is non-portable, as a future release of GNU m4 may change.

• System V m4 supports multiple arguments to defn, and POSIX requires it. This is not yet implemented in GNU m4. Unfortunately, this means it is not possible to mix builtins and other text into a single macro; a helper macro is required.

• POSIX requires an application to exit with non-zero status if it wrote an error message to stderr. This has not yet been consistently implemented for the various builtins that are required to issue an error (such as include (see Section 8.1 [Include], page 47) when a file is unreadable, eval (see Section 11.2 [Eval], page 61) when an argument cannot be parsed, or using m4exit (see Section 13.3 [M4exit], page 70) with a non-numeric argument).

• Some traditional implementations only allow reading standard input once, but GNU m4 correctly handles multiple instances of ‘-’ on the command line.

• POSIX requires m4wrap (see Section 7.5 [M4wrap], page 44) to act in FIFO (first-in, first-out) order, but GNU m4 currently uses LIFO order. Furthermore, POSIX states that only the first argument to m4wrap is saved for later evaluation, bug GNU m4 saves and processes all arguments, with output separated by spaces. However, it is possible to emulate POSIX behavior by including the file ‘examples/ wrapfifo.m4’ from the distribution:

     undivert(`wrapfifo.m4')dnl

     ⇒dnl Redefine m4wrap to have FIFO semantics.
     ⇒define(`_m4wrap_level', `0')dnl
     ⇒define(`m4wrap',
     ⇒`ifdef(`m4wrap'_m4wrap_level,
     ⇒  `define(`m4wrap'_m4wrap_level,
     ⇒  defn(`m4wrap'_m4wrap_level)`$1')',
     ⇒  `builtin(`m4wrap', `define(`_m4wrap_level',
     ⇒  incr(_m4wrap_level))dnl
     ⇒m4wrap'_m4wrap_level)dnl
     ⇒define(`m4wrap'_m4wrap_level, `$1')')')dnl
     include(`wrapfifo.m4')

⇒

     m4wrap(`a`'m4wrap(`c
     ', `d')')m4wrap(`b')

⇒

     ^D
     ⇒abc

• POSIX requires that all builtins that require arguments, but are called without arguments, behave as though empty strings had been passed. For example, a`'define`'b would expand to ab. But GNU m4 ignores certain builtins if they have missing arguments, giving adefineb for the above example.

• Traditional implementations handle define(`f',`1') (see Section 4.1 [Define], page 17) by undefining the entire stack of previous definitions, and if doing undefine(`f') first. GNU m4 replaces just the top definition on the stack, as if doing

popdef(`f') followed by pushdef(`f',`1').

• POSIX requires syscmd (see Section 12.2 [Syscmd], page 66) to evaluate command output for macro expansion, but this appears to be a mistake in POSIX since traditional implementations did not do this. GNU m4 follows traditional behavior in syscmd, and provides the extension esyscmd that provides the POSIX semantics.

• POSIX requires maketemp (see Section 12.5 [Maketemp], page 68) to replace the trailing ‘X’ characters with the m4 process id, giving the same result on identical input, without creating any files, which leaves the door open for a data race in which other processes can create a file by the same name. GNU m4 actually creates a temporary file for each invocation of maketemp, which means that the output of the macro is different even if the input is identical.

• POSIX requires changequote(arg) (see Section 7.2 [Changequote], page 37) to use newline as the close quote, but GNU m4 uses ‘'’ as the close quote. Meanwhile, some traditional implementations use arg as the close quote, making it impossible to nest quotes. For predictable results, never call changequote with just one argument.

• Some implementations of m4 give macros a higher precedence than comments when parsing, meaning that if the start delimiter given to changecom (see Section 7.3 [Changecom], page 40) starts with a macro name, comments are effectively disabled. POSIX does not specify what the precedence is, so the GNU m4 parser recognizes comments, then macros, then quoted strings.

• Traditional implementations allow argument collection, but not string and comment processing, to span file boundaries. Thus, if ‘a.m4’ contains ‘len(’, and ‘b.m4’ contains ‘abc)’, m4 a.m4 b.m4 outputs ‘3’ with traditional m4, but gives an error message that the end of file was encountered inside a macro with GNU m4. On the other hand, traditional implementations do end of file processing for files included with include or sinclude (see Section 8.1 [Include], page 47), while GNU m4 seamlessly integrates the content of those files. Thus include(`a.m4')include(`b.m4') will output ‘3’ instead of giving an error.

• Traditional m4 treats traceon (see Section 6.2 [Trace], page 31) without arguments as a global variable, independent of named macro tracing. Also, once a macro is undefined, named tracing of that macro is lost. On the other hand, when GNU m4 encounters traceon without arguments, it turns tracing on for all existing definitions at the time, but does not trace future definitions; traceoff without arguments turns tracing off for all definitions regardless of whether they were also traced by name; and tracing by name, such as with ‘-tfoo’ at the command line or traceon(`foo') in the input, is an attribute that is preserved even if the macro is currently undefined.

• POSIX requires eval (see Section 11.2 [Eval], page 61) to treat all operators with the same precedence as C. However, GNU m4 currently follows the traditional precedence of other m4 implementations, where bitwise and logical negation (‘~’ and ‘!’) have lower precedence than equality operators, rather than equal precedence with other unary operators. Use explicit parentheses to ensure proper precedence. As extensions to POSIX, GNU m4 treats the shift operators ‘<<’ and ‘>>’ as well-defined on signed integers (even though they are not in C), and adds the exponentiation operator ‘**’.

• POSIX requires translit (see Section 10.5 [Translit], page 57) to treat each character of the second and third arguments literally, but GNU m4 treats ‘-’ as a range operator.

• POSIX requires m4 to honor the locale environment variables of LANG, LC_ALL, LC_ CTYPE, LC_MESSAGES, and NLSPATH, but this has not yet been implemented in GNU

m4.

15.3 Other incompatibilities

There are a few other incompatibilities between this implementation of m4, and the System V version.

• GNU m4 implements sync lines differently from System V m4, when text is being diverted. GNU m4 outputs the sync lines when the text is being diverted, and System V m4 when the diverted text is being brought back. The problem is which lines and file names should be attached to text that is being, or has been, diverted. System V m4 regards all the diverted text as being generated by the source line containing the undivert call, whereas GNU m4 regards the diverted text as being generated at the time it is diverted. The sync line option is used mostly when using m4 as a front end to a compiler. If a diverted line causes a compiler error, the error messages should most probably refer to the place where the diversion were made, and not where it was inserted again.

• GNU m4 makes no attempt at prohibiting self-referential definitions like:

     define(`x', `x')

⇒

     define(`x', `x ')

⇒ There is nothing inherently wrong with defining ‘x’ to return ‘x’. The wrong thing is to expand ‘x’ unquoted. In m4, one might use macros to hold strings, as we do for variables in other programming languages, further checking them with:

     ifelse(defn(`holder'), `value', ...)

In cases like this one, an interdiction for a macro to hold its own name would be a useless limitation. Of course, this leaves more rope for the GNU m4 user to hang himself! Rescanning hangs may be avoided through careful programming, a little like for endless loops in traditional programming languages.

82 GNU M4 macro processor

16 Correct version of some examples

Some of the examples in this manuals are buggy, for demonstration purposes. Correctly working macros are presented here.

The exch macro (see Section 4.2 [Arguments], page 18) as presented requires clients to double quote their arguments. A nicer definition, which lets clients follow the rule of thumb of one level of quoting per level of parentheses, involves adding quotes in the definition of exch, as follows:

     define(`exch', ``$2', `$1'')

⇒

     define(exch(`expansion text', `macro'))

⇒

     macro
     ⇒expansion text

The cleardivert macro (see Section 9.4 [Cleardiv], page 52) cannot, as it stands, be called without arguments to clear all pending diversions. That is because using undivert with an empty string for an argument is different than using it with no arguments at all. Compare the earlier definition with one that takes the number of arguments into account:

     define(`cleardivert',
       `pushdef(`_n', divnum)divert(`-1')undivert($@)divert(_n)popdef(`_n')')

⇒

     divert(`1')one
     divert

⇒

     cleardivert

⇒

     undivert
     ⇒one

⇒

     define(`cleardivert',
       `pushdef(`_num', divnum)divert(`-1')ifelse(`$#', `0',
         `undivert`'', `undivert($@)')divert(_num)popdef(`_num')')

⇒

     divert(`2')two
     divert

⇒

     cleardivert

⇒

     undivert

⇒ The fatal_error macro (see Section 13.3 [M4exit], page 70) does not quite match the

format of internal error messages when invoked inside wrapped text, due to the current limitations of __file__(see Section 13.2 [Location], page 69) when invoked inside m4wrap. Since m4 omits the file and line number from its warning messages when there is no current file (or equivalently, when the current line is 0, since all files start at line 1), a better implementation would be:

define(`fatal_error',
  `errprint(__program__:ifelse(__line__, `0', `',
    `__file__:__line__:')` fatal error: $*
')m4exit(`1')')

⇒

m4wrap(`divnum(`demo of internal message')
fatal_error(`inside wrapped text')')

⇒

^D
 error m4: Warning: excess arguments to builtin `divnum' ignored
⇒0
 error m4: fatal error: inside wrapped text

Appendix A How to make copies of this manual

A.1 GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 51 Franklin St, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.2
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

92 GNU M4 macro processor Appendix B: Indices of concepts and macros 93

Appendix B Indices of concepts and macros

B.1 Index for many concepts

A

arguments to macros . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Arguments to macros . . . . . . . . . . . . . . . . . . . . . . . . . . 18 arguments to macros, special . . . . . . . . . . . . . . . . . . 19 arguments, quoted macro . . . . . . . . . . . . . . . . . . . . . . 15 arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

B

builtins, indirect call of . . . . . . . . . . . . . . . . . . . . . . . . 25

C

call of builtins, indirect . . . . . . . . . . . . . . . . . . . . . . . . 25 call of macros, indirect. . . . . . . . . . . . . . . . . . . . . . . . . 24 changing comment delimiters . . . . . . . . . . . . . . . . . . 40 changing the quote delimiters . . . . . . . . . . . . . . . . . . 37 characters, translating . . . . . . . . . . . . . . . . . . . . . . . . . 57 command line, file names on the . . . . . . . . . . . . . . . . 8 command line, macro definitions on the . . . . . . . . . 5 command line, options . . . . . . . . . . . . . . . . . . . . . . . . . . 4 commands, exit status from shell . . . . . . . . . . . . . . 67 commands, running shell. . . . . . . . . . . . . . . . . . . . . . . 65 commands, running UNIX . . . . . . . . . . . . . . . . . . . . . 65 comment delimiters, changing. . . . . . . . . . . . . . . . . . 40 comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 comments, copied to output . . . . . . . . . . . . . . . . . . . 40 comparing strings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 conditionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 controlling debugging output . . . . . . . . . . . . . . . . . . 33 counting loops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

D

debugging output, controlling . . . . . . . . . . . . . . . . . . 33 debugging output, saving . . . . . . . . . . . . . . . . . . . . . . 34 decrement operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 defining new macros . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 definitions, displaying macro . . . . . . . . . . . . . . . . . . . 31 deleting macros. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 deleting whitespace in input . . . . . . . . . . . . . . . . . . . 37 discarding diverted text. . . . . . . . . . . . . . . . . . . . . . . . 52 displaying macro definitions. . . . . . . . . . . . . . . . . . . . 31 diversion numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 diverted text, discarding . . . . . . . . . . . . . . . . . . . . . . . 52 diverting output to files. . . . . . . . . . . . . . . . . . . . . . . . 49 dumping into frozen file. . . . . . . . . . . . . . . . . . . . . . . . 73

E

error messages, printing. . . . . . . . . . . . . . . . . . . . . . . . 69 evaluation, of integer expressions. . . . . . . . . . . . . . . 61 executing shell commands . . . . . . . . . . . . . . . . . . . . . 65 executing UNIX commands . . . . . . . . . . . . . . . . . . . . 65 exit status from shell commands . . . . . . . . . . . . . . . 67 exiting from m4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 expansion of macros . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 expansion, tracing macro . . . . . . . . . . . . . . . . . . . . . . 31 expressions, evaluation of integer. . . . . . . . . . . . . . . 61 extracting substrings . . . . . . . . . . . . . . . . . . . . . . . . . . 56

F

fast loading of frozen files . . . . . . . . . . . . . . . . . . . . . . 73 FDL, GNU Free Documentation License . . . . . . . 85 file format, frozen file . . . . . . . . . . . . . . . . . . . . . . . . . . 74 file inclusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 51 file names, on the command line . . . . . . . . . . . . . . . . 8 files, diverting output to . . . . . . . . . . . . . . . . . . . . . . . 49 files, names of temporary . . . . . . . . . . . . . . . . . . . . . . 68 for loops. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 formatted output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 frozen file format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 frozen files for fast loading . . . . . . . . . . . . . . . . . . . . . 73

G

GNU extensions . . 17, 18, 24, 25, 34, 48, 51, 55, 57, 59, 66, 73, 77

I

included files, search path for . . . . . . . . . . . . . . . . . . 48 inclusion, of files . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 51 increment operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 indirect call of builtins . . . . . . . . . . . . . . . . . . . . . . . . . 25 indirect call of macros . . . . . . . . . . . . . . . . . . . . . . . . . 24 initialization, frozen states . . . . . . . . . . . . . . . . . . . . . 73 input tokens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 input, saving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 integer arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 integer expression evaluation . . . . . . . . . . . . . . . . . . . 61

L

length of strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 lexical structure of words . . . . . . . . . . . . . . . . . . . . . . 42 License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 local variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

94 GNU M4 macro processor

loops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 loops, counting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

M

macro definitions, on the command line . . . . . . . . . 5 macro expansion, tracing . . . . . . . . . . . . . . . . . . . . . . 31 macro invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 macros, arguments to . . . . . . . . . . . . . . . . . . . . . . 15, 18 macros, displaying definitions . . . . . . . . . . . . . . . . . . 31 macros, expansion of. . . . . . . . . . . . . . . . . . . . . . . . . . . 16 macros, how to define new . . . . . . . . . . . . . . . . . . . . . 17 macros, how to delete. . . . . . . . . . . . . . . . . . . . . . . . . . 21 macros, how to rename . . . . . . . . . . . . . . . . . . . . . . . . 21 macros, indirect call of. . . . . . . . . . . . . . . . . . . . . . . . . 24 macros, quoted arguments to . . . . . . . . . . . . . . . . . . 15 macros, recursive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 macros, special arguments to . . . . . . . . . . . . . . . . . . 19 macros, temporary redefinition of . . . . . . . . . . . . . . 23 messages, printing error. . . . . . . . . . . . . . . . . . . . . . . . 69 multibranches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

N

names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

O

options, command line . . . . . . . . . . . . . . . . . . . . . . . . . . 4 output, diverting to files . . . . . . . . . . . . . . . . . . . . . . . 49 output, formatted . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 output, saving debugging . . . . . . . . . . . . . . . . . . . . . . 34

P

pattern substitution . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 platform macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 platform macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 POSIXLY_CORRECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 printing error messages . . . . . . . . . . . . . . . . . . . . . . . . 69

Q

quote delimiters, changing the . . . . . . . . . . . . . . . . . 37 quoted macro arguments. . . . . . . . . . . . . . . . . . . . . . . 15 quoted string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

R

recursive macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 redefinition of macros, temporary . . . . . . . . . . . . . . 23 regular expressions . . . . . . . . . . . . . . . . . . . . . . . . . 55, 57 reloading a frozen file . . . . . . . . . . . . . . . . . . . . . . . . . . 73 renaming macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 running shell commands . . . . . . . . . . . . . . . . . . . . . . . 65 running UNIX commands. . . . . . . . . . . . . . . . . . . . . . 65

S

saving debugging output . . . . . . . . . . . . . . . . . . . . . . . 34 saving input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 search path for included files . . . . . . . . . . . . . . . . . . . 48 shell commands, exit status from . . . . . . . . . . . . . . 67 shell commands, running. . . . . . . . . . . . . . . . . . . . . . . 65 special arguments to macros . . . . . . . . . . . . . . . . . . . 19 status of shell commands . . . . . . . . . . . . . . . . . . . . . . 67 status, setting m4 exit. . . . . . . . . . . . . . . . . . . . . . . . . . 70 strings, length of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 substitution by regular expression. . . . . . . . . . . . . . 57 substrings, extracting . . . . . . . . . . . . . . . . . . . . . . . . . . 56

T

temporary file names . . . . . . . . . . . . . . . . . . . . . . . . . . 68 temporary redefinition of macros . . . . . . . . . . . . . . . 23 tokens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 tracing macro expansion . . . . . . . . . . . . . . . . . . . . . . . 31 translating characters. . . . . . . . . . . . . . . . . . . . . . . . . . 57

U

undefining macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 UNIX commands, exit status from . . . . . . . . . . . . . 67 UNIX commands, running . . . . . . . . . . . . . . . . . . . . . 65

V

variables, local. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

W

words, lexical structure of. . . . . . . . . . . . . . . . . . . . . . 42

B.2 Index for all m4 macros

References are exclusively to the places where a builtin is introduced the first time.

Appendix B: Indices of concepts and macros 95

__file__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 __gnu__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 __line__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 __os2__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 __program__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 __unix__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 __windows__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

B

builtin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

C

capitalize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 changecom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 changequote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 changeword. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 cleardivert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

D

debugfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 debugmode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 decr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 defn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 divert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 divnum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 dnl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 downcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 dumpdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

E

errprint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 esyscmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

F

fatal_error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 forloop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

I

ifdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ifelse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 incr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 indir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

L

len. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

M

m4exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 m4wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 maketemp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

O

os2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

P

patsubst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 popdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 pushdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

R

regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

S

shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 sinclude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 substr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 syscmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 sysval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

T

traceoff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 traceon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 translit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

U

undefine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 undivert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 unix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 upcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

W

windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

96 GNU M4 macro processor